### v2

# TiM4Rec: An Efficient Sequential Recommendation Model Based on Time-Aware Structured State Space Duality Model

Abstract.
Sequential recommendation represents a pivotal branch of recommendation systems, centered around dynamically analyzing the sequential dependencies between user preferences and their interactive behaviors. Despite the Transformer architecture-based models achieving commendable performance within this domain, their quadratic computational complexity relative to the sequence dimension impedes efficient modeling. In response, the innovative Mamba architecture, characterized by linear computational complexity, has emerged. Mamba4Rec further pioneers the application of Mamba in sequential recommendation. Nonetheless, Mamba 1’s hardware-aware algorithm struggles to efficiently leverage modern matrix computational units, which lead to the proposal of the improved State Space Duality (SSD), also known as Mamba 2. While the SSD4Rec successfully adapts the SSD architecture for sequential recommendation, showing promising results in high-dimensional contexts, it suffers significant performance drops in low-dimensional scenarios crucial for pure ID sequential recommendation tasks. Addressing this challenge, we propose a novel sequential recommendation backbone model, TiM4Rec, which ameliorates the low-dimensional performance loss of the SSD architecture while preserving its computational efficiency. Drawing inspiration from TiSASRec, we develop a time-aware enhancement method tailored for the linear computation demands of the SSD architecture, thereby enhancing its adaptability and achieving state-of-the-art (SOTA) performance in both low and high-dimensional modeling. The code for our model is publicly accessible at https://github.com/AlwaysFHao/TiM4Rec.
Sequential Recommendation, State Space Model(SSM), State Space Duality(SSD), Time-aware Mask Matric
*   The corresponding author.
[…]
journalyear: 2024ccs: Information systems Recommender systems
1. Introduction
Sequential recommendation represents a predictive technology that leverages user’s historical interaction sequence. Unlike traditional recommendation methods—such as collaborative filtering and content based filtering—which rely heavily on static user features and historical behavior, it emphasizes capturing dynamic user behavior characteristic, which is more adept at accommodating rapid shifts in user interests and diverse consumption patterns.
The landscape of sequential recommendation models has evolved significantly, transitioning from early models based on Markov chains (fpmc; hrm) to those leveraging CNNs (tang2018personalized), RNNs (li2017neural; GRU4Rec), GNNs (srgnn; chang2021sequential), notable examples include Caser (tang2018personalized) and GRU4Rec (GRU4Rec). However, these models often suffer from performance limitations inherent to their core architectures, resulting in suboptimal prediction accuracy. The introduction of the Transformer (vaswani2017attention) architecture, grounded in an attention mechanism, has marked a paradigm shift due to its superior performance and parallelizable training capabilities and made it a mainstream framework across various domains that has also significantly influenced the sequential recommendation domain. In recent years, as evidenced by models like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec) effectively adapting the success of Transformers to this area.
Despite their advancements, these early Transformer-based models did not sufficiently account for the unique data characteristics, notably the temporal shifts in user interests inherent to sequential recommendation. Subsequently, researchers introduced enhanced models with time-awareness capabilities. Models such as TiSASRec (TiSASRec) and BTMT (BTMT) extend SASRec by incorporating temporal position encoding, generating additional position encodings based on timestamps of interactions with various items, thus augmenting the expressive power of the attention score matrix. Furthermore, the TiCoSeRec (TiCoSeRec) model employs data augmentation techniques by inserting or masking items within the original interaction sequence, and its approach aims to mitigate the presence of non-uniform sub-sequences and highlight genuine interest points of users. The essence of the aforementioned methods addresses the dynamic nature of users’ interests over time. For instance, as demonstrated in Figure 1, we extract a single user’s interaction record from a real dataset, revealing significant shifts in their interests over time—an essential characteristic of user interaction behavior. Drawing inspiration from this observation, we propose a novel temporal perception enhancement method that offers substantial efficiency improvements over prior temporal enhancement methods, particularly those dominated by TiSASRec.
Figure 1. User interest transfer. Over time, a noticeable shift in users’ points of interest is observed, reflecting a common characteristic of user interaction behaviors.
In contrast to the quadratic computational complexity typically associated with Transformers in the sequence dimension, the Mamba architecture offers the advantage of achieving comparable performance with linear computational complexity.With the Mamba architecture’s advantages being validated in other fields (zhu2024vision; vim-UCAS; xu2024survey; lieber2024jamba), recent studies in the realm of sequential recommendation have increasingly concentrated on applying the Mamba architecture to sequential recommendation systems (mamba4rec; bidirectionalgatedmamba4rec; echomamba4re; ssd4rec). The Mamba framework is divided into two generations: Mamba1 (gu2023mamba), which is based on SSM, and Mamba2 (SSD), which is founded on SSD. A detailed discussion of SSD and SSM is provided in Section 2.
The Mamba4Rec (mamba4rec) model is the Mamba framework’s first attempt at sequential recommendation tasks, which uses the mamba1 architecture to replace the attention mechanism module without compromising performance and significantly improving computational efficiency. Then, the SSD4Rec (ssd4rec) model is developed based on Mamba2 after Mamba4Rec. Due to the enhanced computational efficiency afforded by the Mamba2 architecture, SSD4Rec achieves superior computational performance in high-dimensional feature spaces. However, it is noteworthy that the Mamba2 architecture exhibits a performance decrement relative to Mamba1 in sequential recommendation tasks. Although SSD4Rec attains performance levels comparable to those of Mamba4Rec in high dimension settings through the implementation of variable sequences and enhancements in bidirectional SSD, since the sequential recommendation task based purely on IDs (identifiers) cannot converge efficiently in high-dimensional spaces, it performs less well than that in low-dimensional environments.
If we solely leverage the advantages of the Mamba2 architecture for high dimension modeling in sequential recommendation tasks, without considering the applicability of high-dimensional processing in pure ID-based sequential recommendation scenarios, the associated computational cost may outweigh its benefits. This is particularly pertinent given that high-dimensional modeling is relatively computationally intensive compared to its low-dimensional counterpart, and purely ID-based sequential recommendation tasks may struggle to converge efficiently in this setting. Thus, addressing the challenge of sustaining performance in low-dimensional feature spaces while leveraging the computational advantages of Mamba2 for modeling sequential recommendation tasks constitutes a critical issue necessitating further investigation.
To address the above issues, this study analyzes the factors contributing to the performance degradation of SSD architecture in comparison to SSM architecture in low-dimensional sequential recommendation tasks. We introduce, for the first time, a time-aware enhancement method tailored for SSD architecture to mitigate this performance loss. This research represents the debut of Mamba architecture in the context of sequential recommendation with respect to time-aware enhancement. It is noteworthy that the time-aware enhancement method proposed in this article preserves the linear computational complexity advantage of the SSD architecture in the sequential dimension. Therefore, in comparison with the SASRec model, which represents the Transformer architecture, and the Mamba4Rec model, which represents the SSM architecture, it retains a significant advantage in terms of model training and inference speed.
We designate our work as TiM4Rec and outline its primary contributions as follows:
  • We conduct a pioneering exploration of a time-aware enhancement method tailored for the linear computational complexity of the Mamba architecture in the realm of sequential recommendation.
  • Through the introduction of our proposed time-aware enhancement method, we develop the TiM4Rec model, which effectively reduces performance loss in the SSD architecture relative to the SSM architecture in low-dimensional spaces, enhancing its suitability for ID modeling tasks and improving training efficiency.
  • Comprehensive experiments across three datasets demonstrate the effectiveness of the TiM4Rec model in low-dimensional spaces, showcasing superior performance compared to SSD4Rec, while maintaining the strengths of the SSD architecture in high-dimensional scenarios.
2. PRELIMINARIES
2.1. Problem Statement
Sequential recommendation methodologies leverage the user’s historical interaction sequences to forecast the subsequent potential interaction item for the user. In particular, tasks that focus solely on predicting the next item in a sequence are referred to as seq2item tasks. This work is also based on the seq2item task for its research. Let $\mathcal{U}=\left\{u_{1},u_{2},\cdots,u_{|\mathcal{U}|}\right\}$ be the set of users and $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{|\mathcal{V}|}\right\}$ be the set of items. For any user $u_{k}$, there exists an interaction sequence $\mathcal{S}_{t}^{u_{k}}=\left\{v_{1}^{u_{k}},v_{2}^{u_{k}},\cdots,v_{t}^{u_{k}}\right\}$ at time step $t$, where $v_{i}^{u_{k}}$ represents the i-th item of user $u_{k}$ interaction. Subsequently, by establishing a learnable embedding table =$\mathcal{E}=\left\{e_{1},e_{2},\cdots,e_{|\mathcal{V}|}\right\}\in\mathbb{R}^{|\mathcal{V}|\times N}$, the item IDs are converted from a sparse vector representation to a dense vector representation, where $e_{j}$ denotes the potential dense vector representation of the item $v_{j}$. We transform the input sequence $\mathcal{S}_{t}^{u_{k}}$ into the historical interaction vector representation $p^{u_{k}}_{t}$ for user $u_{k}$ through the mapping $\mathcal{F}$. Subsequently, we compute the inner product between $p^{u_{k}}_{t}$ and all item embedding vectors to obtain the predicted score ranking $r_{t}^{u_{k}}$ for the next interaction item.
$p_{t}^{u_{k}}$ $=\mathcal{F}\left(S_{t}^{u_{k}}\right)$
$r_{t}^{u_{k}}$ $=p_{t}^{u_{k}}\cdot\mathcal{E}$
Finally, we return the IDs of the top $m$ items that the user is likely to interact with, based on the predicted score ranking $r_{t}^{u_{k}}$.
2.2. State Space Models
The State Space Models(SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps the input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
(1) $h^{\prime}(t)$ $=\bm{A}h(t)+\bm{B}x(t),$
$y(t)$ $=\bm{C}h(t)+\bm{D}x(t),$
where $\bm{A}\in\mathbb{R}^{N\times N}$ and $\bm{B},\bm{C}\in\mathbb{R}^{N\times D}$ are learnable matrices. To enable SSM to effectively represent discrete data, it is imperative to discretize the data in accordance with the given step size $\Delta$. An effective discretization method for Equation 1 is the Zero Order Hold (ZOH) method (gu2021combining). Assuming the previous time step is $t-1$ and the current time step is $t$, Formula 1 can be solved by using the constant variation method:
[…]
$=e^{\Delta\bm{A}}h(t-1)+\left(\Delta\bm{A}\right)^{-1}\left(e^{\Delta\bm{A}}-I\right)\Delta\bm{B}x\left(t\right)$
We assume $\overline{\bm{A}}=\exp(\Delta\bm{A})$ and $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$, then the following result can be derived:
(2) $h_{t}$ $=\overline{\bm{A}}h_{t-1}+\overline{\bm{B}}x_{t},$
$y_{t}$ $=\bm{C}h_{t}+\bm{D}x_{t},$
The Structured State Space Model (S4) (gu2021efficiently) is developed from the vanilla SSM. By imposing a structure on the state matrix $\bm{A}\in\mathbb{R}^{D\times N}$ through HiPPO initialization (gu2020hippo), it further enhances the modeling of long-range dependencies.
Based on the foundation established by S4 model, Mamba (gu2023mamba) advances the concept of selectivity by integrating parameter matrices $\Delta\in\mathbb{R}^{T\times D}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, which vary over time. Through these methodological advancements, Mamba is capable of selectively transforming the input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ into the output sequence $\bm{Y}\in\mathbb{R}^{T\times D}$ as delineated by the following equation:
$\bm{Y}_{i}=\sum_{j}\left[\bm{C}_{i}^{\top}\left(\overline{\bm{A}}_{i}\cdot\overline{\bm{B}}_{j}\right)\right]\cdot X_{j}$
$\bm{Y}_{i}=\sum_{j}\left\{\bm{C}_{i}^{\top}\left[\left(\prod_{k=1}^{i-j}e^{\Delta_{k}^{\top}\bm{A}}\right)\cdot\left(\Delta_{j}^{\top}\bm{A}\right)^{-1}\left(e^{\Delta_{j}^{\top}\bm{A}}-\bm{I}\right)\cdot\left(\Delta_{j}^{\top}\bm{B}_{j}\right)\right]\right\}\cdot X_{j}$
Furthermore, the model enhances computational efficiency by employing parallel computation strategies facilitated through the implementation of the Hillis-Steele scan algorithm (hillis1986data). Through the aforementioned enhancements, Mamba demonstrates the capability to attain performance levels that are nearly on par with those of Transformers, while incurring only linear computational complexity with respect to the sequence dimension.
2.3. State Space Duality
The SSM framework, despite employing hardware-aware selective scan algorithms to facilitate parallel computing, encounters limitations in effectively harnessing matrix computation units within the hardware. This inefficacy stems from its non-reliance on matrix operations as a foundational approach. Furthermore, Mamba’s hardware-aware algorithm computational core resides entirely within the GPU’s SRAM, which is typically of limited capacity. This restriction constrains the feature dimensionality in model computations, thereby impeding the ability of the model to operate within higher-dimensional spaces.
From a computational efficiency standpoint, Gu et al. (SSD) innovatively transformed matrix $\bm{A}$ into a scalar value to facilitate SSM matrixization calculation and discovered that this variant of SSM exhibits duality with masked attention mechanisms. To preserve the linear complexity benefit of SSM in the sequence dimension, they leveraged the block properties of semi-separable matrices to implement linear attention computation, culminating in what they termed the Structured State Space Duality.
In essence, SSD, the structure they proposed, initially formulated SSM employing a matrix computation framework:
$\bm{Y}$ $=SSM(\bm{A},\bm{B},\bm{C})(X)=\bm{M}\bm{X}$
$\bm{A}_{m:n}^{\times}$ $\coloneqq\prod_{k=m}^{n}\bm{A}_{k}=\bm{A}_{n}\bm{A}_{n-1}\cdots\bm{A}_{m}$
(3) $\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{1:i}^{\times}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{2:i}^{\times}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{B}_{j}\end{bmatrix}$
In this discussion, we assume the condition where matrices $\bm{A}$ and $\bm{B}$ are not subjected to discretization, allowing us to assert that matrix $\bm{M}_{ij}=\bm{C}_{i}^{\top}\bm{A}_{i-j:i}^{\times}\bm{B}_{j}$. The aforementioned equation delineates the transformation process of the input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ into the output sequence $\bm{Y}\in\mathbb{R}^{T\times D}$ facilitated by the parameter matrices $\bm{A}\in\mathbb{R}^{T\times N}$, $\bm{C}\in\mathbb{R}^{T\times N}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{M}\in\mathbb{R}^{T\times T}$. The inefficiency of SSM in executing matrix operations stems from the computational demands constrained by the recursive computational load of matrix $\bm{A}_{m:n}^{\times}$ in a temporal sequence. Consequently, by simplifying matrix $\bm{A}_{m:n}^{\times}$ to a scalar and leveraging the properties of semi-separable matrices, we can significantly enhance the efficiency of matrix computations involving matrix $\bm{M}$. Under the assumption that $\bm{A}=[1,a_{1},\dots,a_{t-1}]\in\mathbb{R}^{T}$, SSD facilitates the computation of matrix $\bm{M}$ by introducing a 1-Semi Separable (1-SS) matrix $\bm{L}\in\mathbb{R}^{T\times T}$ defined as follows:
(4) $\bm{L}=\begin{bmatrix}1\\ a_{1}&1\\ a_{2}a_{1}&a_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ a_{t-1}\ldots a_{1}&a_{t-1}\ldots a_{2}&\cdots&a_{t-1}&1\end{bmatrix}$
$\bm{M}_{i+1,j+1}=\bm{C}_{i}^{\top}\bm{A}_{i:j}\bm{B}_{j}\coloneqq\bm{C}_{i}^{\top}a_{i}\ldots a_{j+1}\bm{B}_{j}$
Thus, we can equivalently derive the following equation:
$\bm{M}=\bm{L}\circ\bm{C}\bm{B}^{\top}$
$\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{A}_{0:0}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{A}_{2:2}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{i:1}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{i:2}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{A}_{i:i}\bm{B}_{i}\end{bmatrix}$
By considering matrix $\bm{C}$ as matrix $\bm{Q}$, matrix $\bm{B}$ as matrix $\bm{K}$ and matrix $\bm{X}$ as matrix $\bm{V}$ within the framework of the attention mechanism and excluding the Softmax computation, a profound conclusion is reached:
$\bm{Y}=\bm{L}\circ\bm{C}\bm{B}^{\top}\bm{X}=\bm{L}\circ\bm{Q}\bm{K}^{\top}\bm{V}$
This concept is referred to as Duality in the SSD framework. By referring to the tensor contraction order in the calculation of linear attention (LinearAttention)—specifically using the associative property of matrix multiplication where $K^{T}V$ is computed first and subsequently left-multiplied by $Q$—the computational complexity is reduced from quadratic to linear in the sequence dimension without considering causal masks, a property that SSD also achieves.
$\bm{Y}=\left(\bm{Q}\bm{K}^{T}\right)\bm{V}\quad\Rightarrow\quad\bm{Y}=\bm{Q}\left(\bm{K}^{T}\bm{V}\right)$
$\bm{Y}=\left(\bm{C}\bm{B}^{T}\right)\bm{X}\quad\Rightarrow\quad\bm{Y}=\bm{C}\left(\bm{B}^{T}\bm{X}\right)$
To integrate causal masking without compromising linear computational complexity, the block properties of the 1-SS matrix $\bm{L}$ can be leveraged in conjunction with segment accumulation. This approach facilitates the implementation of causal linear attention, referred to as Structured Masked Attention (SMA) (SSD) in SSD. Since the SSD architecture is derived from the Mamba model, models based on the SSD architecture are referred to as Mamba2. Although SSD offers significant advantages in terms of computational load—particularly in tasks with high feature dimensions—the performance of SSD in sequential recommendation tasks is inferior compared to Mamba1. We provide a detailed analysis of this issue in the following sections.
3. METHOD
Figure 2. The overview of TiM4Rec. The overall framework of TiM4Rec, which is based on time-aware enhancement, enhances the performance of the SSD architecture in the domain of sequential recommendation by effectively processing the time differences in interactions between items.
In this section, we offer a detailed introduction to our proposed framework, TiM4Rec. Following an overview of the overall structure of TiM4Rec, we delve into the reasons behind the performance degradation of SSD compared to SSM in low-dimensional scenarios within sequential recommendation tasks. Subsequently, we explain how our time-aware enhancement method, utilizing our proposed time-aware masked matrix with linear computational complexity, compensates for this loss.
3.1. Overview of TiM4Rec
As shown in Figure 2, the overall logic of TiM4Rec adheres to the description in section 2.1, However, an additional user interaction unix timestamp sequence $\mathcal{T}^{u_{k}}=\left\{t_{1},t_{2},\cdots,t_{L}\right\}\in\mathbb{R}^{L}$ is incorporated into the input data, with $\mathcal{T}^{u_{k}}$ corresponding one to one with the user interation sequence $\mathcal{S}^{u_{k}}$, where $L$ representing interaction sequence length. The sequence $\mathcal{S}^{u_{k}}$ can be transformed into an input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ composed of dense vectors by the embedding table $\mathcal{E}$. Subsequently, a displacement subtraction is performed on sequence $\mathcal{T}^{u_{k}}$, and a zero is appended to the begining to generate the intearaction time difference sequence $\mathcal{D}^{u_{k}}=\left\{d_{1},d_{2},\cdots,d_{L}\right\}\in\mathbb{R}^{L}$, which be used in later stages. The Time-aware SSD Block enhances the model’s ability to capture potential user interests by integrating sequence $\mathcal{D}^{u_{k}}$ into a discretized SSD kernel. For the sake of convenience in the subsequent discussion, the superscript $u_{k}$ be omitted.
3.2. Time Difference Processing
As mentioned above, for the input interaction timestamp sequence $\mathcal{T}$, we transform it into an interaction time difference sequence $\mathcal{D}$ using the following method:
(5) $\mathcal{D}_{i}=\begin{cases}0,\quad&i=1\\ \mathcal{T}_{i}-\mathcal{T}_{i-1},\quad&i\in\left[2,L\right]\end{cases}$
Subsequently, we apply dropout (srivastava2014dropout) and layer normalization (ba2016layer) to sequence $\mathcal{D}$ in order to eliminate data noise and optimize the data distribution for subsequent analysis.
3.3. Analyzing the Masked Matrix in SSD
As discussed in section 2.3, the critical advancement from SSM to SSD resides in the scalarization of the transition matrix $\bm{A}$. When considering this modification from an attention perspective, the central aspect involves the alterations in the mask matric $\bm{L}$. Although the recursive operational constraints of matrix $\bm{A}$ in SSM prevent the direct and precise formulation of the masked matrix within eq. 3, it remains feasible to examine an equivalent expression for the final masked matrix:
$\bm{L}_{\text{SSM}}=\begin{bmatrix}I\\ \bm{A}_{1}&I\\ \bm{A}_{2}\bm{A}_{1}&\bm{A}_{2}&I\\ \vdots&\vdots&\ddots&\ddots\\ \bm{A}_{1:i}^{\times}&\bm{A}_{2:i}^{\times}&\cdots&\bm{A}_{i-1:i}^{\times}&I\end{bmatrix}$
As illustrated in figure 3, in contrast to the masked matrix of SSD presented in eq. 4, the fundamental change in the masked matrix $L_{\text{SSM}}$ is that its mask coefficients are applied in a point-to-point manner on the individual item $\bm{E}_{i}=\left[e_{1},e_{2},\cdots,e_{n}\right]\in\mathbb{R}^{N}$ (The $\bm{A}$ matrix follows its definitions in SSM and SSD, respectively.):
[…]
In instances where the $\bm{E}$ is in low dimension, its relative contribution to information is diminished, necessitating a more refined filtering of $\bm{E}$ by the masked matrix. Consequently, the point-to-point masking employed in SSM demonstrates superior performance in low-dimensional contexts. Notably, from the perspective of masking strategies, SSD exhibits a greater inclination towards the Transformer architecture than SSM.
$\bm{L}_{MSA}=\begin{bmatrix}1\\ 1&1\\ 1&1&1\\ \vdots&\vdots&\ddots&\ddots\\ 1&1&\cdots&1&1\end{bmatrix},$
MSA stands for Vanilla Mask Self Attention Mechanism
However, owing to the semi-separable matrix structure inherent in the masked matrix of SSD, there is a pronounced attenuation effect on the interaction items that appear earlier in the sequence. This feature aligns SSD more closely with the actual user interaction patterns observed in real-world recommendation systems. Consequently, it is anticipated that, under equivalent conditions in low-dimensional settings, the SSM architecture exhibit superior performance in the realm of sequential recommendation, followed by SSD, with the Transformer architecture trailing behind. This hypothesis is further corroborated by the findings presented in our subsequent experimental section.
Figure 3. The distinction in mask formulation between SSD and SSM is noteworthy. For an individual item, SSM can be viewed as an equivalent point-to-point mask, whereas SSD resembles the global mask utilized in Transformer architectures.
Observing the differences between the equivalent mask matrices of SSD and SSM, it can be seen that the core of their mask matrices is recursively generated based on the state transition matrix $\bm{A}$. As described in section 2.3, the only difference is that SSD achieves more efficient matrix multiplication calculations by scalar quantization of matrix $\bm{A}$. Therefore, addressing the loss of the masked matrix caused by the scalar quantization of matrix $\bm{A}$ while maintaining the efficient matrix multiplication calculation of SSD has become the key to improvement.
3.4. Time-aware Masked Matrix
Let’s reconsider how the time difference matrix should be expressed in the context of attention mechanisms. The method proposed by TiSASRec (TiSASRec) is centered around acquiring the time position encoding for each item in the interaction sequence through interaction timestamps, ultimately leading to the generation of the corresponding attention bias matrix based on these encodings. Given that the Mamba architecture does not depend on positional encoding—as substantiated by the findings in Mamba4Rec (mamba4rec) and SSD4Rec (ssd4rec)—it is essential to directly explore the equivalent representation of the TiSASRec approach. Specifically, we should investigate how the resulting attention bias matrix encapsulates the temporal differences in product interactions.
For a sequence of interactive timestamps $\mathcal{T}$, we can derive the relative time difference matrix $\hat{\mathcal{D}}$ between each product through the following transformation:
(6) $\hat{\mathcal{D}}=\begin{bmatrix}\hat{d}_{11}\\ \mathbf{\underline{\hat{d}_{21}}}&\hat{d}_{22}\\ \hat{d}_{31}&\mathbf{\underline{\hat{d}_{32}}}&\hat{d}_{33}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{d}_{L1}&\hat{d}_{L2}&\cdots&\mathbf{\underline{\hat{d}_{LL-1}}}&\hat{d}_{LL}\end{bmatrix},\quad\hat{d}_{mn}=t_{m}-t_{n}$
Directly applying matrix $\hat{\mathcal{D}}$ to the masked matrix $\bm{L}$ (as described in eq. 4) compromises the semi-separable matrix property of $\bm{L}$. This compromise prevents us from leveraging the efficient matrix computation techniques offered by the SSD architecture, consequently resulting in a return to quadratic computational complexity. Therefore, it is imperative to modify the insertion method of matrix $\hat{\mathcal{D}}$.
When we construct a new matrix $\mathcal{D}$ by selecting only the values of the underlined elements in matrix $\hat{\mathcal{D}}$, we effectively adopt the form described in eq. 5. This approach allows us to integrate matrix $\mathcal{D}$ into the masked matrix $\bm{L}$ with linear complexity. Initially, we apply matrix $\mathcal{D}$ to the scalarization matrix $\bm{A}$, which results in the corresponding positions in the masked matrix $\bm{L}$ containing the relevant time difference information (Due to the presence of layer normalization, the actual value of $d_{1}$ will be non-zero.):
$\hat{\bm{A}}_{i}=\hat{a}_{i}=\mathcal{D}_{i}\cdot\bm{A}_{i}=\hat{d}_{ii-1}\cdot a_{i}=d_{i}\cdot a_{i}$
$\hat{\bm{A}}_{m:n}\coloneqq\hat{a}_{n}\hat{a}_{n-1}\cdots\hat{a}_{m}=a_{n}d_{n}\cdot a_{n-1}d_{n-1}\cdots a_{m}d_{m}$
(7) $\bm{L}=\begin{bmatrix}1\\ \hat{a}_{1}&1\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&1\end{bmatrix}$
This can be regarded as an equivalent approximation of $d_{ij}$ as follows:
$\hat{d}_{ij}\approx\hat{d}_{ii-1}\hat{d}_{i-1i-2}\cdots\hat{d}_{j+1j}=d_{i}d_{i-1}\cdots d_{j}$
In the context of SSD kernel computations, multiplication is effectively converted into accumulation operations via the logarithm function. This transformation offers advantageous support for the aforementioned approximation of $\hat{d}_{ij}$.
$\log\left(d_{i}d_{i-1}\cdots d_{j}\right)=\log\left(d_{i}\right)+\log\left(d_{i-1}\right)+\cdots+\log\left(d_{j}\right)$
Through the approximation method mentioned above, we accomplish the time-aware enhancement of the SSD architecture. This is achieved without disrupting the semi-separable property of the masked matrix $\bm{L}$.
3.5. Time-aware SSD Block
Despite providing a detailed derivation of the SSD in Section 2.3, the actual code implementation of the SSD kernel retains the discretization process. Therefore, further analysis is required. Notably, in Section 2.2, we derive the discretization for the SSM, but it is important to highlight that the actual computational process is approximated in both SSM and SSD implementations. Specifically, for matrix $\overline{\bm{A}}$, the implementation approximates the exponential operation by directly computing $\overline{\bm{A}}\approx\Delta\bm{A}$. Meanwhile, for matrix $\overline{\bm{B}}$, given the Maclaurin series expansion of $e^{x}$ as $1+x+\frac{x^{2}}{2}+\cdots$, the following approximation can be made:
$\overline{\bm{B}}$ $\approx\left(\Delta\bm{A}\right)^{-1}\left(\bm{I}+\Delta\bm{A}-\bm{I}\right)\cdot\Delta\bm{B}$
$=\left(\Delta\bm{A}\right)^{-1}\Delta\bm{A}\cdot\Delta\bm{B}$
$=\Delta\bm{B}$
In this approximation, we establish the following definitions: the state transition coefficient $\bm{A}\in\mathbb{R}^{1}$. By inputting $\bm{X}\in\mathbb{R}^{T\times D}$, we derive matrices $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, along with the discretization parameter $\Delta\in\mathbb{R}^{T}$:
$\bm{X},\bm{B},\bm{C},\Delta=\bm{X}W+b$
Where the weight matrix $W\in\mathbb{R}^{D\times(D+2N+1)}$ and the bias matrix $b\in\mathbb{R}^{D+2N+1}$. Subsequently, a causal convolution transformation is applied to the matrices $\bm{X}$, $\bm{B}$ and $\bm{C}$:
$\bm{X}_{t},\bm{B}_{t},\bm{C}_{t}$ $=\sigma\left[(\bm{X}_{t},\bm{B}_{t},\bm{C}_{t})^{\top}*\omega\right]$
$\text{where}\quad\mathcal{Q}_{t}=\mathcal{P}_{t}*\omega$ $\coloneqq\sum_{m=0}^{K-1}\mathcal{P}_{\max\left(t-m,0\right)}\cdot\omega_{m}$
Where the $\omega\in\mathbb{R}^{K}$ is convolution kernel and the $\sigma$ represents non-linear activation functions.
Before proceeding with the derivation, it is prudent to examine how the time difference vector $\mathcal{D}$ is integrated into the discretized SSD. Regarding the time difference matrix $\mathcal{D}$, we note that the implicit interest conveyed by the same interaction time difference can vary across different user interaction contexts. Therefore, we initially apply a time-varying transformation to the time difference matrix using the following transformation:
$\alpha^{\mathcal{D}}=\sigma\left[\sigma\left(\mathcal{D}W_{1}^{\mathcal{D}}+b_{1}^{\mathcal{D}}\right)W_{2}^{\mathcal{D}}+b_{2}^{\mathcal{D}}\right]$
$\mathcal{D}=\alpha^{\mathcal{D}}\cdot\mathcal{D}$
Where the weight matrix $W_{1}^{\mathcal{D}},W_{2}^{\mathcal{D}}\in\mathbb{R}^{T\times T}$ and the bias matrix $b_{1}^{\mathcal{D}},b_{2}^{\mathcal{D}}\in\mathbb{R}^{T}$. Considering the convergence of interaction time differences within a short time frame, we apply causal convolution to the time difference matrix $\mathcal{D}$ to enhance the implicit features of user interest points.
$\mathcal{D}_{t}=\sigma\left(\mathcal{D}_{t}*\omega^{\mathcal{D}}\right)\coloneqq\sigma\left(\sum_{m=0}^{K-1}\mathcal{D}_{\max\left(t-m,0\right)}\cdot\omega^{\mathcal{D}}_{m}\right)$
Next, we integrate the enhanced time difference matrix $\mathcal{D}$ with the discrete parameter $\Delta$, and obtain matrices $\overline{\bm{A}}$ and $\overline{\bm{B}}$ through the following transformation:
$\hat{\Delta}=\Delta\cdot\mathcal{D},\quad\overline{\bm{A}}=\hat{\bm{A}}=\hat{\Delta}\cdot\bm{A},\quad\overline{\bm{B}}=\hat{\Delta}\cdot\bm{B}$
Finally, the following equation can be derived to map the input $\bm{X}$ to the output $\bm{Y}\in\mathbb{R}^{T\times D}$($\bm{L}$ refers to eq. 7.):
$\bm{Y}=\bm{L}\circ\bm{C}\overline{\bm{B}}^{\top}\bm{X}=\bm{L}\circ\bm{C}\bm{B}^{\top}\left(\hat{\Delta}^{\top}\bm{X}\right)$
Due to the preservation of the semi-separable matrix property of $\bm{L}$, the matrix computation acceleration method discussed in Section 2.3 remains applicable.
To extract deeper user interest features, multiple layers of Time-aware SSD blocks are stacked. Consequently, after the Time-aware SSD, the features are transformed by adding a Feed Forward Network layer to adapt to the next semantic space. It is noteworthy that to adapt the time difference vector $\mathcal{D}$ to the feature semantic space of the next layer, gate residual (he2016deep) processing is applied to the input time difference vector $\mathcal{D}$ for the subsequent layer.
3.6. Prediction Layer
By utilizing the Time-aware SSD Block, input $\bm{X}$ can be transformed into a user interest feature sequence $\bm{P}\in\mathbb{R}^{L\times D}$. The last element $p_{t}$ of the feature sequence $\bm{P}$ is then extracted as the interest feature corresponding to the current user interaction sequence.
$\hat{\mathcal{Y}}=Softmax\left(p_{t}\cdot\mathcal{E}\right)$
As stated in Section 2.1, the predicted scores $\hat{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{V}|}$ for all items are obtained by calculating the inner product of $p_{t}$ and the embedding table of all items. However, a notable difference is that the actual score calculation incorporates the Softmax function, even though the Softmax operation does not influence the ranking of the scores.
3.7. Complexity Analysis
Models based on the Vanilla Transformer, such as SASRec, require a total of $O(L^{2}N)$ floating-point operations (FLOPs). In contrast, thanks to the computational improvements provided by Mamba, models based on SSM and SSD, such as Mamba4Rec and SSD4Rec, only require $O(LN^{2})$ FLOPs. In practical application scenarios of sequential recommendation, including news and product recommendations, it is often the case that $L\gg N$. Thus, sequential recommendation models based on Mamba demonstrate superior advantages. TiSASRec enhances its performance by integrating time-aware capabilities via time difference encoding. Despite the theoretical FLOPS remaining in the order of $O(L^{2}N)$, the practical computational and memory demands exceed those of SASRec by more than four times. This substantial increase is attributed to the inherent computational constraints associated with time difference encoding. Consequently, while TiSASRec demonstrates improved time sensitivity, it necessitates a significantly higher computational overhead and memory allocation compared to SASRec, posing potential challenges in scalability and efficiency. In contrast, TiM4Rec preserves the benefits of the SSD architecture by employing a time difference masked matrix characterized by linear computational complexity. This approach maintains the computational efficiency, demanding only $O(LN^{2})$ FLOPs, and incurs only a modest increase in additional computational overhead. Notably, the generation of the time-aware masked matrix in TiM4Rec is solely based on scalar computations. This characteristic ensures that the additional computational complexity does not escalate in high-dimensional modeling scenarios, thereby preserving the superior matrix operation efficiency inherent to the SSD4Rec architecture.
Table 1. Comparison of computational complexity between representative sequential recommendation methods, where ”M.M.” refers to ”Matrix Multiplication”.
Architecture FLOPs Memory M.M.
SASRec Attention $L^{2}N$ $L^{2}N$
Mamba4Rec SSM $LN^{2}$ $LN^{2}$
SSD4Rec SSD $LN^{2}$ $LN$
TiSASRec Attention $L^{2}N$ $L^{2}N$
TiM4Rec SSD $LN^{2}$ $LN$
4. Experiments
Table 2. Sequential recommendation performance on the ML-1M, Beauty, and KuaiRand datasets. Here, R@k refers to Recall@K, N@K denotes NDCG@K, and M@K represents MRR@K. The best and second-best results are indicated in bold and underlined, respectively. ”Improv.” signifies the performance improvement of TiM4Rec compared to the second-best results. The asterisk in SSD4Rec* denotes that this is our own replicated version.
Model Type & Model $\rightarrow$ CNN RNN Transformer SSM SSD Improv.
Dataset Metric Caser GRU4Rec SASRec BERT4Rec TiSASRec Mamba4Rec SSD4Rec* TiM4Rec
MovieLens-1M R@10 0.2156 0.2985 0.3060 0.2800 0.3147 0.3253 0.3199 0.3310 1.75%
R@20 0.3083 0.3937 0.4050 0.3853 0.4250 0.4354 0.4240 0.4338 -0.30%
R@50 0.4515 0.5397 0.5455 0.5377 0.5608 0.5707 0.5649 0.5770 1.10%
N@10 0.1157 0.1700 0.1754 0.1584 0.1834 0.1891 0.1841 0.1932 2.17%
N@20 0.1391 0.1941 0.2005 0.1852 0.2106 0.2169 0.2104 0.2194 1.15%
N@50 0.1675 0.2232 0.2285 0.2154 0.2376 0.2436 0.2384 0.2477 1.68%
M@10 0.0853 0.1307 0.1357 0.1214 0.1424 0.1474 0.1425 0.1512 2.58%
M@20 0.0917 0.1373 0.1425 0.1288 0.1498 0.1551 0.1498 0.1584 2.13%
M@50 0.0963 0.1420 0.1471 0.1337 0.1542 0.1593 0.1542 0.1629 2.26%
Amazon Beauty R@10 0.0402 0.0563 0.0851 0.0352 0.0802 0.0838 0.0806 0.0854 0.35%
R@20 0.0618 0.0856 0.1194 0.0563 0.1147 0.1185 0.1146 0.1204 0.84%
R@50 0.1090 0.1376 0.1759 0.0999 0.1734 0.1802 0.1710 0.1800 -0.11%
N@10 0.0199 0.0297 0.0425 0.0172 0.0405 0.0435 0.0423 0.0446 2.53%
N@20 0.0253 0.0371 0.0511 0.0225 0.0491 0.0522 0.0509 0.0533 2.11%
N@50 0.0346 0.0473 0.0623 0.0311 0.0605 0.0644 0.0620 0.0651 1.09%
M@10 0.0138 0.0217 0.0294 0.0117 0.0282 0.0311 0.0306 0.0321 3.22%
M@20 0.0153 0.0237 0.0317 0.0132 0.0306 0.0335 0.0329 0.0345 2.99%
M@50 0.0167 0.0253 0.0335 0.0145 0.0325 0.0354 0.0347 0.0363 2.54%
Kuai Rand R@10 0.0801 0.1020 0.1055 0.0938 0.1057 0.1094 0.1055 0.1109 1.37%
R@20 0.1344 0.1659 0.1704 0.1537 0.1710 0.1768 0.1717 0.1774 0.34%
R@50 0.2561 0.3017 0.3074 0.2873 0.3060 0.3154 0.3088 0.3202 1.52%
N@10 0.0395 0.0564 0.0584 0.0510 0.0590 0.0608 0.5880 0.0611 0.49%
N@20 0.0531 0.0724 0.0747 0.0660 0.0753 0.0777 0.0754 0.0779 0.26%
N@50 0.0770 0.0911 0.1016 0.0923 0.1019 0.1050 0.1024 0.1060 0.95%
M@10 0.0273 0.0428 0.0443 0.0382 0.0450 0.0461 0.0449 0.0463 0.43%
M@20 0.0310 0.0471 0.0487 0.0422 0.0494 0.0508 0.0494 0.0508 -
M@50 0.0347 0.0513 0.0529 0.0464 0.0536 0.0552 0.0536 0.0552 -
4.1. Experimental Setup
4.1.1. Datasets
The performance of our proposed model is evaluated through experiments conducted on three publicly available datasets, which have previously been utilized as evaluation benchmarks in several classical sequential recommendation models.
  • MovieLens-1M (harper2015movielens): A dataset containing approximately 1 million user ratings for movies, collected from the MovieLens platform.
  • Amazon-Beauty (mcauley2015image): The user review dataset collected in the beauty category on the Amazon platform was compiled up to the year 2014.
  • Kuai Rand (kuairand): Acquired from the recommendation logs of the application Kuaishou, the dataset includes millions of interactions involving items that were randomly displayed.
For each user, we sort their interaction records according to the timestamps, thereby generating an interaction sequence for each individual. And then retain only those users and items that are associated with a minimum of five interaction records, in accordance with the methodology established in prior research (kang2018self). The statistics of these datasets are shown in Table 3.
4.1.2. Baseline
To demonstrate the superiority of our proposed method, we select a set of representative sequential recommendation models to serve as baselines. Notably, even though SSD4Rec (ssd4rec) has not been released as open-source, we implemented a variant termed SSD4Rec*, adhering to the traditional sequential recommendation framework to validate the efficacy of our time-aware approach. This implementation is strictly based on the SSD architecture and does not incorporate features such as variable-length sequences and bidirectional SSD methods proposed in SSD4Rec. The exclusion of these features is intentional, as they are generally regarded as universal enhancement techniques, particularly in the context of handling variable-length sequences.
  • Caser (tang2018personalized): A classical sequential recommendation model based on CNN.
  • GRU4Rec (GRU4Rec): A RNN-based method constructed by Gated Recurrent Units (GRU).
  • SASRec (kang2018self): The first exploration model of applying Transformer in the field of sequential recommendation.
  • BERT4Rec (sun2019bert4rec): A bidirectional attention sequential recommendation model following the BERT (bert) model paradigm.
  • TiSASRec (TiSASRec): A time-aware enhanced sequential recommendation model based on SASRec: Considered as paradigm in time-aware sequential recommendation research.
  • Mamba4Rec (mamba4rec): A pioneering model that explores the application of the Mamba architecture in the domain of sequential recommendation.
  • SSD4Rec (ssd4rec): The pioneering sequential recommendation model leveraging the SSD architecture, exploits its inherent advantages over the SSM architecture for high-dimensional modeling, and integrates variable-length sequence training techniques to enhance model performance.
4.1.3. Evaluation Metrics
We employ well-established recommendation performance metrics including Hit Ratio (HR@K), Normalized Discounted Cumulative Gain (NDCG@K), and Mean Reciprocal Rank (MRR@K) as the criteria for experimental assessment. In calculating these metrics, it is necessary to extract the top K scores from the model’s final predicted ranking. We choose K values of 10, 20, and 50 to encompass short, medium, and long prediction lengths.
Table 3. Statistics of the experimented datasets.
Dataset ML-1M Beauty KuaiRand
#Users 6,040 22,363 23,951
#Items 3,416 12,101 7,111
#Interactions 999,611 198,502 1,134,420
Avg.Length 165.5 8.9 47.4
Max.Length 2,314 389 809
Sparsity 95.15% 99.93% 99.33%
4.1.4. Implementation Details
Our experimental procedures adhere to the paradigm requirements set by the PyTorch (paszke2019pytorch) and RecBole (recbole[1.2.0]). All models in the experiment are implemented with a model dimension of 64, which is the conventional choice for sequential recommendation models based on pure ID modeling. For models based on the Mamba architecture, such as Mamba4Rec, SSD4Rec* and TiM4Rec, the SSM state factor was consistently set to 32, the kernel size for 1D causal convolution is 4, and the block expansion factor for linear projections is 2. It is notable that the 1D causal convolutional layer within the Mamba architecture constrains the multi-head parameter to multiples of 8. However, in the context of sequential recommendation, this parameter typically holds values of 2 or 4. Consequently, we modify the 1D causal convolutional layer to accommodate the specific modeling tasks of sequential recommendation, ultimately standardizing it to a value of 4. The backbone layers across all models are consistently configured to a depth of 2. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M and KuaiRand. The Adam optimizer (kingma2014adam) is used uniformly with a learning rate set to 0.01. A batch size of 2048 is employed for training, and the validated batch size is 4096. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens1M and 50 for Amazon-Beauty and KuaiRand datasets. We adhere to the optimal parameter configurations specific to each baseline, ensuring that the core parameters remain consistent across all models, and make adjustments to maximize performance to the extent feasible.
4.2. Performance Comparison
We initially conduct a comparative analysis of the recommendation performance between our proposed model and several baseline models across three benchmark datasets. The results of this comparison are presented in Table 2. Through a detailed examination of the experimental outcomes, we are able to derive the following conclusions:
  • In low-dimensional settings, the SSD architecture experiences a considerable decline in recommendation performance compared to the SSM architecture. However, when compared with models founded on Transformer architectures, it exhibits enhanced performance. This observation aligns with our theoretical analysis presented in Section 3.3.
  • TiM4Rec exhibits a significant performance enhancement when compared to SSD4Rec*, which is modeled on a purely SSD-based architecture, thereby confirming the efficacy of the Time-aware SSD Block. Furthermore, TiM4Rec demonstrates a slight performance advantage over Mamba4Rec. This observation underscores the effectiveness of leveraging temporal differences in user interactions to capture shifts in user interests, effectively mitigating the performance decline observed when transitioning from the SSM architecture to the SSD architecture.
4.3. Dimension comparison
To assess the suitability of modeling pure ID sequence recommendation tasks in lower dimensions, we compare the experimental results presented in the SSD4Rec paper at 256 dimensions. These results are shown in Table 4. Our research findings clearly demonstrate that the Mamba4Rec model tends to exhibit superior performance in 64-dimensional space. Additionally, the SSD4Rec* model that we replicate, which does not utilize variable-length sequences or bidirectional SSDs, showed only a slight decline in performance at 64 dimensions compared to the results presented in the original SSD4Rec paper at 256 dimensions. Notably, the performance of SSD4Rec at 256 dimensions is even inferior to that Mamba4Rec at 64 dimensions. These findings provide robust evidence that pure ID sequence recommendation tasks are more effectively modeled in lower-dimensional spaces. Moreover, the computational overhead associated with 256 dimensions is substantially higher, underscoring the advantages of utilizing our model. In addtional, a key advantage of the SSD architecture over the SSM architecture is its superior capability in high-dimensional modeling. To validate this advantage, we conduct a comparative analysis of the performance of our proposed method, TiM4Rec, at a dimensionality of 256. The results confirm that TiM4Rec effectively preserves the benefits of the SSD architecture even in high-dimensional settings.
Table 4. The performance of sequential recommendation models based on SSM and SSD is compared on the ML1M dataset in 64 and 256 dimensions. The definitions are aligned with the provisions of Table 2.
Dimension $\rightarrow$ 64D 256D
Metric Mamba4Rec SSD4Rec* TiM4Rec Mamba4Rec SSD4Rec TiM4Rec
R@10 0.3253 0.3199 0.3310 0.3124 0.3152 0.3270
R@20 0.4354 0.4240 0.4338 0.4103 0.4194 0.4272
N@10 0.1891 0.1841 0.1932 0.1847 0.1889 0.1938
N@20 0.2169 0.2104 0.2194 0.2094 0.2145 0.2191
M@10 0.1474 0.1425 0.1512 0.1456 0.1495 0.1529
M@20 0.1551 0.1498 0.1584 0.1523 0.1565 0.1598
Figure 4. Regarding the performance radar chart of Table 4
4.4. Efficiency Analysis
In this section, we verify whether TiM4Rec can retain the inherent advantages of SSD architecture during model training and inference. Table 5 presents the training and inference times for representative models built on Transformer and Mamba architectures, all executed on a single NVIDIA RTX 3090 GPU.
Observing the experimental results, we can draw the following conclusions:
  • In both high-dimensional and low-dimensional spaces, TiM4Rec effectively preserves the inherent computational efficiency of the SSD architecture, necessitating only a minimal additional computational load when compared to SSD4Rec*, which is exclusively based on the SSD architecture. Notably, when compared with SASRec and Mamba4Rec, TiM4Rec achieves substantial enhancements in computational efficiency, attributed to leveraging the high-efficiency computing capabilities of the SSD architecture. This demonstrates TiM4Rec’s ability to significantly optimize computational processes while minimizing overhead, thereby underscoring its advantage in scenarios demanding efficient and scalable solutions.
  • Although TiSASRec maintains the computational complexity akin to that of SASRec, its time-aware methodology introduces notable inefficiencies, leading to a substantial escalation in actual computational load. By contrast, TiM4Rec demonstrates remarkable efficiency by necessitating only minimal additional computational resources across both high-dimensional and low-dimensional domains. The experimental results robustly validate the efficacy of the proposed time-aware enhancement method, which maintains linear computational complexity, underscoring its potential as an effective solution for efficient sequential recommendation systems.
Table 5. Comparison of training and inference times across various models using the ML-1M dataset. All models have a dimensionality of 64, except those with 256d suffix.
Model Architecture Training Time Inference Time
SASRec Attention 207.42s 0.81s
TiSASRec Attention 1149.57s 1.89s
Mamba4Rec SSM 111.76s 0.26s
SSD4Rec* SSD 74.43s 0.18s
TiM4Rec SSD 83.57s 0.21s
SSD4Rec*-256d SSD 338.23s 0.64s
TiM4Rec-256d SSD 342.38s 0.64s
It is noteworthy that although we only compare the computational efficiency of our own replicated version of SSD4Rec*, the actual computational efficiency of SSD4Rec* surpasses that of SSD4Rec because SSD4Rec* eliminates the bidirectional SSD processing and the variable length sequence methods originally proposed in SSD4Rec, which come with additional overhead.
4.5. Ablation Study
To validate the efficacy of our proposed time-aware algorithm along with other critical design elements, we perform a series of ablation studies using the ML-1M dataset. The procedure for the systematic elimination of each component is detailed as follows:
  • w/o Time : Remove the proposed time-aware algorithm from the SSD module, rendering the model equivalent to our replication of the SSD4Rec* model mentioned earlier.
  • w/o FFN : To assess the necessity of the FFN layer’s nonlinear transformation ability in extracting sequential features, we eliminated this layer as part of our ablation study.
  • N layer : Performance of the Model with N Stacked Layers of Time-aware SSD Blocks.
Table 6. Results of ablation studies on the ML-1M dataset
Component R@10 R@20 N@10 N@20 M@10 M@20
Full 0.3310 0.4338 0.1932 0.2194 0.1512 0.1584
w/o Time 0.3199 0.4240 0.1841 0.2104 0.1425 0.1498
w/o FFN 0.3247 0.4278 0.1842 0.2103 0.1415 0.1487
1 layer 0.3088 0.4146 0.1773 0.2040 0.1373 0.1446
3 layer 0.3310 0.4323 0.1916 0.2172 0.1490 0.1560
Based on the results presented in Table 6, it is evident that the time-aware enhancement method our propose significantly improves performance. The ablation experiment findings further underscore the critical role of the FFN layer’s feature mapping ability in the extraction of sequence features.
5. Conclusion and Future Work
In conclusion, this study makes significant strides in the advancement of sequential recommendation systems by thoroughly analyzing the SSM and SSD architectures. Our enhancements to the SSD framework, particularly through the integration of a time-aware mask matrix, effectively address the challenges associated with performance degradation in low-dimensional spaces. This novel approach not only preserves but also enhances the computational efficiency of SSDs, thereby setting a new benchmark in the field. Moreover, our research marks a pioneering effort in exploring temporal enhancement techniques within the Mamba series architecture, achieving remarkable state-of-the-art performance.
Our contributions offer valuable insights into the development of efficient and robust recommendation systems, opening up new avenues for further exploration. Notably, our work emphasizes the importance of temporal dynamics in improving recommendation accuracy and efficiency. Moving forward, we are excited to extend our research into the domain of multi-modal sequential recommendation. By doing so, we aim to develop algorithms that are not only more aligned with the intricate demands of real-world recommendation systems but also capable of leveraging diverse data modalities for improved prediction and recommendation outcomes. This future direction holds promise for broadening the applicability and effectiveness of sequential recommendation algorithms, ultimately leading to more personalized and context-aware user experiences.
Acknowledgements.
The authors declare no potential conflict of interest. We extend our gratitude to Master Guodong Sun from Northwestern Polytechnical University for his valuable revision suggestions for this article. The work was supported by Humanity and Social Science Foundation of Ministry of Education of China (No.18YJA630037, 21YJA630054), Zhejiang Province Soft Science Research Program Project (No.2024C350470).

### v3

@@
 Abstract.
-Sequential recommendation represents a pivotal branch of recommendation systems, centered around dynamically analyzing the sequential dependencies between user preferences and their interactive behaviors.
-Despite the Transformer architecture-based models achieving commendable performance within this domain, their quadratic computational complexity relative to the sequence dimension impedes efficient modeling.
-In response, the innovative Mamba architecture, characterized by linear computational complexity, has emerged.
-Mamba4Rec further pioneers the application of Mamba in sequential recommendation.
-Nonetheless, Mamba 1’s hardware-aware algorithm struggles to efficiently leverage modern matrix computational units, which lead to the proposal of the improved State Space Duality (SSD), also known as Mamba 2.
-While the SSD4Rec successfully adapts the SSD architecture for sequential recommendation, showing promising results in high-dimensional contexts, it suffers significant performance drops in low-dimensional scenarios crucial for pure ID sequential recommendation tasks.
-Addressing this challenge, we propose a novel sequential recommendation backbone model, TiM4Rec, which ameliorates the low-dimensional performance loss of the SSD architecture while preserving its computational efficiency.
-Drawing inspiration from TiSASRec, we develop a time-aware enhancement method tailored for the linear computation demands of the SSD architecture, thereby enhancing its adaptability and achieving state-of-the-art (SOTA) performance in both low and high-dimensional modeling.
-The code for our model is publicly accessible at https://github.com/AlwaysFHao/TiM4Rec.
-Sequential Recommendation, State Space Model(SSM), State Space Duality(SSD), Time-aware Mask Matric
+The Sequential Recommendation modeling paradigm is shifting from Transformer to Mamba architecture, which comprises two generations: Mamba1, based on the State Space Model (SSM), and Mamba2, based on State Space Duality (SSD).
+Although SSD offers superior computational efficiency compared to SSM, it suffers performance degradation in sequential recommendation tasks, especially in low-dimensional scenarios that are critical for these tasks.
+Considering that time-aware enhancement methods are commonly employed to mitigate performance loss, our analysis reveals that the performance decline of SSD can similarly be fundamentally compensated by leveraging mechanisms in time-aware methods.
+Thus, we propose integrating time-awareness into the SSD framework to address these performance issues.
+However, integrating current time-aware methods, modeled after TiSASRec, into SSD faces the following challenges: 1) the complexity of integrating these transformer-based mechanisms with the SSD architecture, and 2) the computational inefficiency caused by the need for dimensionality expansion of time-difference modeling.
+To overcome these challenges, we introduce a novel Time-aware Structured Masked Matrix that efficiently incorporates time-aware capabilities into SSD.
+Building on this, we propose Time-Aware Mamba for Recommendation (TiM4Rec), which mitigates performance degradation in low-dimensional SSD contexts while preserving computational efficiency.
+This marks the inaugural application of a time-aware enhancement method specifically tailored for the Mamba architecture within the domain of sequential recommendation.
+Extensive experiments conducted on three real-world datasets demonstrate the superiority of our approach.
+The code for our model is accessible at https://github.com/AlwaysFHao/TiM4Rec.
+Sequential Recommendation, State Space Model(SSM), State Space Duality(SSD), Mamba, Time-awareness
 *   The corresponding author.
@@
 1. Introduction
-Sequential recommendation represents a predictive technology that leverages user’s historical interaction sequence.
-Unlike traditional recommendation methods—such as collaborative filtering and content based filtering—which rely heavily on static user features and historical behavior, it emphasizes capturing dynamic user behavior characteristic, which is more adept at accommodating rapid shifts in user interests and diverse consumption patterns.
-The landscape of sequential recommendation models has evolved significantly, transitioning from early models based on Markov chains (fpmc; hrm) to those leveraging CNNs (tang2018personalized), RNNs (li2017neural; GRU4Rec), GNNs (srgnn; chang2021sequential), notable examples include Caser (tang2018personalized) and GRU4Rec (GRU4Rec).
+Sequential Recommendation (SR) is an important branch of Recommendation Systems (RS), leveraging the user’s historical interaction sequence to predict the items that the user may interact with next (RSSurvey2005; SRSurvey2018).
+Unlike traditional methods—such as collaborative filtering (CollaborativeFiltering2001) and content-based filtering (VBPR2016)—that predominantly depend on static user features and past behaviors, SR focuses on capturing dynamic user behavior patterns, making it more effective in adapting to rapid shifts in user interests and diverse consumption habits (Deezer2022; SASRec2018; RSSurvey2005).
+The landscape of SR models has evolved significantly, transitioning from early models based on Markov chains (HRM2016; FPMC2010) to those leveraging Convolutional Neural Networks (CNNs) (TangW2018), Recurrent Neural Networks (RNNs) (GRU4Rec2016; LiRCRLM2017), Graph Neural Networks (GNNs) (ChangGZHNSJ02021; SRGNN2019), notable examples include Caser (TangW2018) and GRU4Rec (GRU4Rec2016).
 However, these models often suffer from performance limitations inherent to their core architectures, resulting in suboptimal prediction accuracy.
-The introduction of the Transformer (vaswani2017attention) architecture, grounded in an attention mechanism, has marked a paradigm shift due to its superior performance and parallelizable training capabilities and made it a mainstream framework across various domains that has also significantly influenced the sequential recommendation domain.
-In recent years, as evidenced by models like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec) effectively adapting the success of Transformers to this area.
-Despite their advancements, these early Transformer-based models did not sufficiently account for the unique data characteristics, notably the temporal shifts in user interests inherent to sequential recommendation.
-Subsequently, researchers introduced enhanced models with time-awareness capabilities.
-Models such as TiSASRec (TiSASRec) and BTMT (BTMT) extend SASRec by incorporating temporal position encoding, generating additional position encodings based on timestamps of interactions with various items, thus augmenting the expressive power of the attention score matrix.
-Furthermore, the TiCoSeRec (TiCoSeRec) model employs data augmentation techniques by inserting or masking items within the original interaction sequence, and its approach aims to mitigate the presence of non-uniform sub-sequences and highlight genuine interest points of users.
-The essence of the aforementioned methods addresses the dynamic nature of users’ interests over time.
-For instance, as demonstrated in Figure 1, we extract a single user’s interaction record from a real dataset, revealing significant shifts in their interests over time—an essential characteristic of user interaction behavior.
-Drawing inspiration from this observation, we propose a novel temporal perception enhancement method that offers substantial efficiency improvements over prior temporal enhancement methods, particularly those dominated by TiSASRec.
-Figure 1. User interest transfer. Over time, a noticeable shift in users’ points of interest is observed, reflecting a common characteristic of user interaction behaviors.
-In contrast to the quadratic computational complexity typically associated with Transformers in the sequence dimension, the Mamba architecture offers the advantage of achieving comparable performance with linear computational complexity.With the Mamba architecture’s advantages being validated in other fields (zhu2024vision; vim-UCAS; xu2024survey; lieber2024jamba), recent studies in the realm of sequential recommendation have increasingly concentrated on applying the Mamba architecture to sequential recommendation systems (mamba4rec; bidirectionalgatedmamba4rec; echomamba4re; ssd4rec).
-The Mamba framework is divided into two generations: Mamba1 (gu2023mamba), which is based on SSM, and Mamba2 (SSD), which is founded on SSD.
-A detailed discussion of SSD and SSM is provided in Section 2.
-The Mamba4Rec (mamba4rec) model is the Mamba framework’s first attempt at sequential recommendation tasks, which uses the mamba1 architecture to replace the attention mechanism module without compromising performance and significantly improving computational efficiency.
-Then, the SSD4Rec (ssd4rec) model is developed based on Mamba2 after Mamba4Rec.
-Due to the enhanced computational efficiency afforded by the Mamba2 architecture, SSD4Rec achieves superior computational performance in high-dimensional feature spaces.
-However, it is noteworthy that the Mamba2 architecture exhibits a performance decrement relative to Mamba1 in sequential recommendation tasks.
-Although SSD4Rec attains performance levels comparable to those of Mamba4Rec in high dimension settings through the implementation of variable sequences and enhancements in bidirectional SSD, since the sequential recommendation task based purely on IDs (identifiers) cannot converge efficiently in high-dimensional spaces, it performs less well than that in low-dimensional environments.
-If we solely leverage the advantages of the Mamba2 architecture for high dimension modeling in sequential recommendation tasks, without considering the applicability of high-dimensional processing in pure ID-based sequential recommendation scenarios, the associated computational cost may outweigh its benefits.
-This is particularly pertinent given that high-dimensional modeling is relatively computationally intensive compared to its low-dimensional counterpart, and purely ID-based sequential recommendation tasks may struggle to converge efficiently in this setting.
-Thus, addressing the challenge of sustaining performance in low-dimensional feature spaces while leveraging the computational advantages of Mamba2 for modeling sequential recommendation tasks constitutes a critical issue necessitating further investigation.
-To address the above issues, this study analyzes the factors contributing to the performance degradation of SSD architecture in comparison to SSM architecture in low-dimensional sequential recommendation tasks.
-We introduce, for the first time, a time-aware enhancement method tailored for SSD architecture to mitigate this performance loss.
-This research represents the debut of Mamba architecture in the context of sequential recommendation with respect to time-aware enhancement.
-It is noteworthy that the time-aware enhancement method proposed in this article preserves the linear computational complexity advantage of the SSD architecture in the sequential dimension.
-Therefore, in comparison with the SASRec model, which represents the Transformer architecture, and the Mamba4Rec model, which represents the SSM architecture, it retains a significant advantage in terms of model training and inference speed.
-We designate our work as TiM4Rec and outline its primary contributions as follows:
-  • We conduct a pioneering exploration of a time-aware enhancement method tailored for the linear computational complexity of the Mamba architecture in the realm of sequential recommendation.
-  • Through the introduction of our proposed time-aware enhancement method, we develop the TiM4Rec model, which effectively reduces performance loss in the SSD architecture relative to the SSM architecture in low-dimensional spaces, enhancing its suitability for ID modeling tasks and improving training efficiency.
-  • Comprehensive experiments across three datasets demonstrate the effectiveness of the TiM4Rec model in low-dimensional spaces, showcasing superior performance compared to SSD4Rec, while maintaining the strengths of the SSD architecture in high-dimensional scenarios.
-2. PRELIMINARIES
-2.1. Problem Statement
-Sequential recommendation methodologies leverage the user’s historical interaction sequences to forecast the subsequent potential interaction item for the user.
-In particular, tasks that focus solely on predicting the next item in a sequence are referred to as seq2item tasks.
-This work is also based on the seq2item task for its research.
-Let $\mathcal{U}=\left\{u_{1},u_{2},\cdots,u_{|\mathcal{U}|}\right\}$ be the set of users and $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{|\mathcal{V}|}\right\}$ be the set of items.
-For any user $u_{k}$, there exists an interaction sequence $\mathcal{S}_{t}^{u_{k}}=\left\{v_{1}^{u_{k}},v_{2}^{u_{k}},\cdots,v_{t}^{u_{k}}\right\}$ at time step $t$, where $v_{i}^{u_{k}}$ represents the i-th item of user $u_{k}$ interaction.
-Subsequently, by establishing a learnable embedding table =$\mathcal{E}=\left\{e_{1},e_{2},\cdots,e_{|\mathcal{V}|}\right\}\in\mathbb{R}^{|\mathcal{V}|\times N}$, the item IDs are converted from a sparse vector representation to a dense vector representation, where $e_{j}$ denotes the potential dense vector representation of the item $v_{j}$.
-We transform the input sequence $\mathcal{S}_{t}^{u_{k}}$ into the historical interaction vector representation $p^{u_{k}}_{t}$ for user $u_{k}$ through the mapping $\mathcal{F}$.
-Subsequently, we compute the inner product between $p^{u_{k}}_{t}$ and all item embedding vectors to obtain the predicted score ranking $r_{t}^{u_{k}}$ for the next interaction item.
-$p_{t}^{u_{k}}$ $=\mathcal{F}\left(S_{t}^{u_{k}}\right)$
-$r_{t}^{u_{k}}$ $=p_{t}^{u_{k}}\cdot\mathcal{E}$
-Finally, we return the IDs of the top $m$ items that the user is likely to interact with, based on the predicted score ranking $r_{t}^{u_{k}}$.
+The introduction of the Transformer architecture (Transformer2017), grounded in an attention mechanism, has marked a paradigm shift due to its superior performance and parallelizable training capabilities and made it a mainstream framework across various domains that have also significantly influenced the sequential recommendation domain (BTMT2024; SASRec2018; BERT4Rec2019; AFMN2022).
+Although Transformer-based sequential recommendation models, such as SASRec (SASRec2018), have demonstrated promising performance, their inherent quadratic computational load of the Transformer architecture poses significant challenges for efficient sequential recommendation modeling (RecBLR2024; Mamba4Rec2024; LRURec2024).
+To mitigate the Transformer’s inherent computational burden, numerous models with linear complexity have been proposed (e.g., HiPPO (HiPPO2020), S4 (S4-2022), LRU (LRU2023) and RWKV (RWKV2023)).
+Among these, Mamba (Mamba2023) strikes a commendable balance between performance and efficiency.
+Mamba is derived from the State Space Model (SSM), building upon prior works (e.g., HiPPO (HiPPO2020), S4 (S4-2022), and S5 (S5-2023)), and has been widely and successfully applied across various domains (MambaNLP2024; VisionMamba2024; SSMEC2024).
+For efficient sequential recommendation modeling, Mamba4Rec (Mamba4Rec2024) is the first to adapt the Mamba architecture to SR.
+Experimental results indicate that the Mamba architecture holds greater potential than the Transformer in SR tasks (RecBLR2024; Mamba4Rec2024).
+Subsequently, Dao et al.
+(SSD2024) optimized the architecture from a parallel computing perspective, proposing a novel architecture called State Space Duality (SSD) based on Mamba, also referred to as Mamba2.
+SSD demonstrates superior efficiency compared to SSM and exhibits enhanced performance in high-dimensional modeling (SSD2024); however, its performance in low-dimensional settings is suboptimal (further analysis will be provided later).
+Although SSD4Rec (SSD4Rec2024) achieves better performance in higher dimensions than Mamba4Rec through techniques such as variable-length sequence training and bidirectional SSD, it does not outperform Mamba4Rec in lower dimensions.
+Pure ID sequential recommendation typically achieves optimal performance in low dimensions (BERT2019; GRU4Rec2016; SASRec2018; LRURec2024), and high-dimensional modeling contradicts the objective of efficient sequential recommendation.
+If we solely leverage the advantages of the Mamba2 architecture for high-dimensional modeling in SR tasks, without considering the applicability of high-dimensional processing in pure ID-based SR scenarios, the associated computational cost may outweigh its benefits.
+Therefore, addressing the performance deficit of the SSD architecture in low dimensions becomes a crucial challenge for efficient sequential recommendation modeling.
+Figure 1. User interest transfer. Over time, a noticeable shift in the user’s points of interest is observed, reflecting a common characteristic of user interaction behaviors.
+User interaction behaviors typically exhibit clear temporal patterns (TiCoSeRec2023; TiSASRec2020).
+As illustrated in Figure 1, we extract an interaction sequence from a real-world dataset (Amazon2014), clearly demonstrating that user interests tend to change systematically over time.
+Previous Transformer-based SR models have not fully leveraged these temporal dynamics (MusicRec2020; TiSASRec2020; MOJITO2023).
+Consequently, numerous time-aware enhancement models have emerged (BTMT2024; TimelyRec2021; MEANTIME2020; TiSASRec2020; MOJITO2023; TAT4SRec2023) based on the TiSASRec (TiSASRec2020) modeling paradigm, aiming to improve the recommendation performance of Transformers by effectively incorporating time-awareness.
+By analyzing the reasons for the performance degradation of SSD compared to SSM, we find that, theoretically, time-aware enhancement methods can address the performance loss issue of SSD in low-dimension.
+Therefore, we propose integrating time-aware enhancement methods into the SSD architecture to tackle this problem.
+However, the existing time-aware augmentation methods encounter the following issues when integrated to SSD architectures:
+Challenge I: Incompatibility with the SSD architecture.
+Current time-aware enhancement methods based on Transformers cannot be effectively integrated into the computational core of SSD.
+Specifically, time-aware enhancement methods following the TiSASRec paradigm fundamentally apply time-difference information to the attention score matrix.
+However, since SSD lacks an explicit representation of the attention score matrix during its computation, existing time-aware enhancement methods are incompatible with SSD’s architecture.
+Challenge II: Low computational efficiency.
+In the paradigm of time-aware enhancement methods based on TiSASRec, it is necessary to increase the dimensionality of the time difference matrix to obtain a time-awareness score matrix, which is then integrated into the attention score matrix.
+This approach results in significant additional computational resource consumption, rendering existing time-aware enhancement methods ineffective for efficient sequential recommendation tasks.
+To address the aforementioned challenges, we introduce, for the first time, a time-aware enhancement method specifically tailored for SSD architectures that maintain computational efficiency while mitigating performance degradation issues.
+Specifically, we propose the Time-aware Structured Masked Matrix, which integrates time-awareness into the SSD architecture.
+This method involves scalar-level transformations along the sequence dimension to convert time difference information into a structured mask matrix suitable for SSD architectures.
+Compared to traditional time-aware enhancement approaches, our method significantly improves computational efficiency.
+It is noteworthy that the time-aware enhancement method proposed in this paper preserves the linear computational complexity advantage of the SSD architecture in the sequence length.
+Therefore, in comparison with the SASRec model (SASRec2018), which represents the Transformer architecture, and the Mamba4Rec model (Mamba4Rec2024), which represents the SSM architecture, it retains a significant advantage in terms of model training and inference speed.
+We designate our work as TiM4Rec(Time-aware Mamba for Recommendation) and outline its primary contributions as follows:
+  • To the best of our knowledge, we conduct a pioneering exploration of a time-aware enhancement method tailored for the linear computational complexity of the Mamba architecture in the field of sequential recommendation.
+  • Our theoretical analysis reveals the underlying reasons for the performance degradation of SSD compared to SSM.
+In response, we propose a targeted time-aware enhancement method designed to mitigate the performance loss of SSD in sequential recommendation tasks.
+  • We integrate time-awareness into the SSD architecture by constructing a Time-aware Structured Masked Matrix.
+In contrast to the time-aware enhancement method modeled on TiSASRec, which requires a computational complexity of $O(T^{2}N)$, our approach achieves efficient time-awareness with only $O(T)$ computational complexity.
+  • Experiments on three real-world public datasets demonstrate the superiority of our proposed TiM4Rec compared with Transformer-based and SSM-based baselines in terms of both effectiveness and efficiency.
+2. RELATED WORK
+2.1. Sequential Recommendation
+Previous research on sequential recommendation systems focused primarily on utilizing Recurrent Neural Networks (RNNs) (GRU4Rec2016; LiRCRLM2017).
+However, due to the performance limitations of RNNs and their inefficiency in parallel computation, an increasing number of successful studies in sequential recommendation have shifted towards employing Transformer-based approaches (BTMT2024; SASRec2018; BERT4Rec2019; AFMN2022).
+The most representative model is SASRec (SASRec2018), which captures user interest features from user interaction sequences through the causal self-attention mechanism (Transformer2017).
+Owing to the powerful modeling capability of self-attention, SASRec achieves strong recommendation performance.
+However, the Transformer architecture requires a quadratic computational load with respect to the sequence length, making it inefficient for sequence modeling (we will analyze the computational complexity in detail in Section 3.6).
+In the application of recommender systems, various platforms typically require real-time recommendations, as excessive latency can detrimentally impact user experience (TencentRec2015; MaNLD2020).
+Consequently, the investigation of efficient sequential recommendation methods has become imperative (RecBLR2024; LRURec2024).
+Recently, many models with linear computational loads have emerged as alternatives to the original Transformer architecture (HiPPO2020; S4-2022; LRU2023; RWKV2023; S5-2023), among which Mamba (Mamba2023) has achieved notable success and has been promoted in various other fields (MambaNLP2024; VisionMamba2024; SSMEC2024).
+Mamba4Rec (Mamba4Rec2024) is the first to adapt the Mamba architecture to the sequential recommendation domain.
+In essence, it employs a modeling method similar to SASRec, merely replacing the self-attention mechanism module in SASRec with the Mamba1 module.
+Since Mamba1 uses state transitions to finely control the flow of information within sequences, rather than the fully connected modeling approach of self-attention that considers each vector’s contribution equally, it better aligns with user interaction habits (i.e., recent interactions more accurately reflect users’ next interaction tendencies).
+Thus, Mamba4Rec achieves better performance while maintaining efficiency.
+With the introduction of the Mamba2 (SSD2024) architecture, SSD4Rec (SSD4Rec2024) has been applied to the sequential recommendation domain for the first time, further enhancing time efficiency compared to Mamba4Rec.
+Although SSD4Rec surpasses Mamba4Rec in high-dimensional modeling through additional processes such as variable-length sequence input and bidirectional modeling, this improvement is based on the premise that Mamba4Rec experiences a performance drop in high-dimensional settings (which we will verify in the experimental section).
+Prioritizing high-dimensional modeling contradicts the original intention of efficient modeling.
+For pure ID sequential recommendation, modeling is inherently more suitable in lower dimensions.
+2.2. Time-aware methods in SR
+As mentioned in the introduction, user interaction behaviors typically exhibit clear temporal patterns (TiCoSeRec2023; TiSASRec2020).
+Consequently, researchers have introduced time-aware enhancement methods into the field of sequential recommendation.
+TiSASRec (TiSASRec2020) represents the first investigation of a time-aware enhanced sequential recommendation algorithm based on the Transformer architecture.
+In essence, TiSASRec enhances the computation of the attention score matrix by constructing a time difference matrix and generating time difference positional encodings based on the difference values.
+MEANTIME (MEANTIME2020) enhances the learning of various temporal encoding patterns by utilizing multiple attention heads.
+TimelyRec (TimelyRec2021) models heterogeneous temporal patterns at different temporal granularities.
+MOJITO (MOJITO2023) captures diverse fine-grained temporal patterns simultaneously by integrating an attention mechanism with Gaussian mixtures.
+TAT4SRec (TAT4SRec2023) employs an encoder-decoder architecture to model timestamps and interaction items separately.
+Although these methods achieve improved performance compared to TiSASRec, they share the same modeling paradigm as TiSASRec and suffer from common limitations such as inefficiency and restriction to attention-based mechanisms.
+The methodological paradigm of TiSASRec, while capable of enhancing the performance of Transformers in sequential recommendation through the integration of time difference matrices, incurs a significant increase in computational overhead.
+This is due to the necessity of embedding the time difference matrices and executing attention computations before they can be incorporated into the final attention score matrix.
+Such an approach contradicts the principle of efficient computation.
+Therefore, discovering a method to directly integrate time difference matrices into the attention score matrix is crucial for applying time-aware enhancement techniques to sequential recommendation modeling efficiently.
 2.3. State Space Models
-The State Space Models(SSM) is a sequence modeling framework based on linear ordinary differential equations.
-It maps the input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
+The State Space Model(SSM) is a sequence modeling framework based on linear ordinary differential equations.
+It maps the input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{D}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
 (1) $h^{\prime}(t)$ $=\bm{A}h(t)+\bm{B}x(t),$
@@
 To enable SSM to effectively represent discrete data, it is imperative to discretize the data in accordance with the given step size $\Delta$.
-An effective discretization method for Equation 1 is the Zero Order Hold (ZOH) method (gu2021combining).
-Assuming the previous time step is $t-1$ and the current time step is $t$, Formula 1 can be solved by using the constant variation method:
+An effective discretization method for Eq.
+1 is the Zero Order Hold (ZOH) method (SSM2021).
+Assuming the previous time step is $t-1$ and the current time step is $t$, Eq.
+1 can be solved by using the constant variation method:
 […]
 […]
-$=e^{\Delta\bm{A}}h(t-1)+\left(\Delta\bm{A}\right)^{-1}\left(e^{\Delta\bm{A}}-I\right)\Delta\bm{B}x\left(t\right)$
+$=\exp(\Delta\bm{A})h(t-1)+\left(\Delta\bm{A}\right)^{-1}\left(\exp(\Delta\bm{A})-I\right)\Delta\bm{B}x\left(t\right),$
 We assume $\overline{\bm{A}}=\exp(\Delta\bm{A})$ and $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$, then the following result can be derived:
@@
 $y_{t}$ $=\bm{C}h_{t}+\bm{D}x_{t},$
-The Structured State Space Model (S4) (gu2021efficiently) is developed from the vanilla SSM.
-By imposing a structure on the state matrix $\bm{A}\in\mathbb{R}^{D\times N}$ through HiPPO initialization (gu2020hippo), it further enhances the modeling of long-range dependencies.
-Based on the foundation established by S4 model, Mamba (gu2023mamba) advances the concept of selectivity by integrating parameter matrices $\Delta\in\mathbb{R}^{T\times D}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, which vary over time.
+The Structured State Space Model (S4) (S4-2022) is developed from the vanilla SSM.
+By imposing a structure on the state matrix $\bm{A}\in\mathbb{R}^{D\times N}$ through HiPPO (HiPPO2020) initialization, it further enhances the modeling of long-range dependencies.
+Based on the foundation established by S4 model, Mamba (Mamba2023) advances the concept of selectivity by integrating parameter matrices $\Delta\in\mathbb{R}^{T\times D}$, $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, which vary over time.
 Through these methodological advancements, Mamba is capable of selectively transforming the input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ into the output sequence $\bm{Y}\in\mathbb{R}^{T\times D}$ as delineated by the following equation:
-$\bm{Y}_{i}=\sum_{j}\left[\bm{C}_{i}^{\top}\left(\overline{\bm{A}}_{i}\cdot\overline{\bm{B}}_{j}\right)\right]\cdot X_{j}$
-$\bm{Y}_{i}=\sum_{j}\left\{\bm{C}_{i}^{\top}\left[\left(\prod_{k=1}^{i-j}e^{\Delta_{k}^{\top}\bm{A}}\right)\cdot\left(\Delta_{j}^{\top}\bm{A}\right)^{-1}\left(e^{\Delta_{j}^{\top}\bm{A}}-\bm{I}\right)\cdot\left(\Delta_{j}^{\top}\bm{B}_{j}\right)\right]\right\}\cdot X_{j}$
-Furthermore, the model enhances computational efficiency by employing parallel computation strategies facilitated through the implementation of the Hillis-Steele scan algorithm (hillis1986data).
-Through the aforementioned enhancements, Mamba demonstrates the capability to attain performance levels that are nearly on par with those of Transformers, while incurring only linear computational complexity with respect to the sequence dimension.
+(3) $\bm{Y}_{i}$ $=\sum_{j}\left[\bm{C}_{i}^{\top}\left(\overline{\bm{A}}_{i}\cdot\overline{\bm{B}}_{j}\right)\right]\cdot X_{j},$
+$\overline{\bm{A}}_{i}$ $=\prod_{k=1}^{i-j}\exp(\Delta_{k}^{\top}\bm{A}),$
+$\overline{\bm{B}}_{j}$ $=\left(\Delta_{j}^{\top}\bm{A}\right)^{-1}\left(\exp(\Delta_{j}^{\top}\bm{A})-\bm{I}\right)\cdot\left(\Delta_{j}^{\top}\bm{B}_{j}\right),$
+Additionally, Mamba leverages CUDA to implement the Parallel Scan algorithm (Parallelizing2018) and draws inspiration from FlashAttention (FlashAttention2022) by employing hardware-aware parallel computation strategies to enhance computational efficiency.
+Through the aforementioned enhancements, Mamba demonstrates the capability to attain performance levels that are nearly on par with those of Transformers, while incurring only linear computational complexity with respect to the sequence length.
 2.4. State Space Duality
@@
 This restriction constrains the feature dimensionality in model computations, thereby impeding the ability of the model to operate within higher-dimensional spaces.
-From a computational efficiency standpoint, Gu et al.
-(SSD) innovatively transformed matrix $\bm{A}$ into a scalar value to facilitate SSM matrixization calculation and discovered that this variant of SSM exhibits duality with masked attention mechanisms.
+From a computational efficiency standpoint, Dao et al.
+(SSD2024) innovatively transformed matrix $\bm{A}$ into a scalar value to facilitate SSM matrixization calculation and discovered that this variant of SSM exhibits duality with masked attention mechanisms.
 To preserve the linear complexity benefit of SSM in the sequence dimension, they leveraged the block properties of semi-separable matrices to implement linear attention computation, culminating in what they termed the Structured State Space Duality.
 In essence, SSD, the structure they proposed, initially formulated SSM employing a matrix computation framework:
-$\bm{Y}$ $=SSM(\bm{A},\bm{B},\bm{C})(X)=\bm{M}\bm{X}$
-$\bm{A}_{m:n}^{\times}$ $\coloneqq\prod_{k=m}^{n}\bm{A}_{k}=\bm{A}_{n}\bm{A}_{n-1}\cdots\bm{A}_{m}$
-(3) $\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{1:i}^{\times}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{2:i}^{\times}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{B}_{j}\end{bmatrix}$
+(4) $\bm{Y}=SSM(\bm{A},\bm{B},\bm{C})(X)=\bm{M}\bm{X},$
+(5) $\bm{A}_{m:n}^{\times}\coloneqq\prod_{k=m}^{n}\bm{A}_{k}=\bm{A}_{n}\bm{A}_{n-1}\cdots\bm{A}_{m},$
+$\bm{M}$ $=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2}\bm{A}_{1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{1:i}^{\times}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{2:i}^{\times}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{B}_{j}\end{bmatrix},$
 In this discussion, we assume the condition where matrices $\bm{A}$ and $\bm{B}$ are not subjected to discretization, allowing us to assert that matrix $\bm{M}_{ij}=\bm{C}_{i}^{\top}\bm{A}_{i-j:i}^{\times}\bm{B}_{j}$.
@@
 Under the assumption that $\bm{A}=[1,a_{1},\dots,a_{t-1}]\in\mathbb{R}^{T}$, SSD facilitates the computation of matrix $\bm{M}$ by introducing a 1-Semi Separable (1-SS) matrix $\bm{L}\in\mathbb{R}^{T\times T}$ defined as follows:
-(4) $\bm{L}=\begin{bmatrix}1\\ a_{1}&1\\ a_{2}a_{1}&a_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ a_{t-1}\ldots a_{1}&a_{t-1}\ldots a_{2}&\cdots&a_{t-1}&1\end{bmatrix}$
-$\bm{M}_{i+1,j+1}=\bm{C}_{i}^{\top}\bm{A}_{i:j}\bm{B}_{j}\coloneqq\bm{C}_{i}^{\top}a_{i}\ldots a_{j+1}\bm{B}_{j}$
+(6) $\bm{L}=\begin{bmatrix}1\\ a_{1}&1\\ a_{2}a_{1}&a_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ a_{t-1}\ldots a_{1}&a_{t-1}\ldots a_{2}&\cdots&a_{t-1}&1\end{bmatrix},$
+$\bm{M}_{i+1,j+1}=\bm{C}_{i}^{\top}\bm{A}_{i:j}\bm{B}_{j}\coloneqq\bm{C}_{i}^{\top}a_{i}\ldots a_{j+1}\bm{B}_{j},$
 Thus, we can equivalently derive the following equation:
-$\bm{M}=\bm{L}\circ\bm{C}\bm{B}^{\top}$
-$\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{A}_{0:0}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{A}_{2:2}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{i:1}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{i:2}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{A}_{i:i}\bm{B}_{i}\end{bmatrix}$
+(7) $\bm{M}=\bm{L}\circ\bm{C}\bm{B}^{\top},$
+$\bm{M}=\begin{bmatrix}\bm{C}_{0}^{\top}\bm{A}_{0:0}\bm{B}_{0}&&&&\\ \bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{0}&\bm{C}_{1}^{\top}\bm{A}_{1:1}\bm{B}_{1}&&&\\ \bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{0}&\bm{C}_{2}^{\top}\bm{A}_{2:1}\bm{B}_{1}&\bm{C}_{2}^{\top}\bm{A}_{2:2}\bm{B}_{2}&&\\ \vdots&\vdots&\ddots\\ \bm{C}_{i}^{\top}\bm{A}_{i:1}\bm{B}_{0}&\bm{C}_{i}^{\top}\bm{A}_{i:2}\bm{B}_{1}&\cdots&\bm{C}_{i}^{\top}\bm{A}_{i:i}\bm{B}_{i}\end{bmatrix}.$
 By considering matrix $\bm{C}$ as matrix $\bm{Q}$, matrix $\bm{B}$ as matrix $\bm{K}$ and matrix $\bm{X}$ as matrix $\bm{V}$ within the framework of the attention mechanism and excluding the Softmax computation, a profound conclusion is reached:
-$\bm{Y}=\bm{L}\circ\bm{C}\bm{B}^{\top}\bm{X}=\bm{L}\circ\bm{Q}\bm{K}^{\top}\bm{V}$
-This concept is referred to as Duality in the SSD framework.
-By referring to the tensor contraction order in the calculation of linear attention (LinearAttention)—specifically using the associative property of matrix multiplication where $K^{T}V$ is computed first and subsequently left-multiplied by $Q$—the computational complexity is reduced from quadratic to linear in the sequence dimension without considering causal masks, a property that SSD also achieves.
-$\bm{Y}=\left(\bm{Q}\bm{K}^{T}\right)\bm{V}\quad\Rightarrow\quad\bm{Y}=\bm{Q}\left(\bm{K}^{T}\bm{V}\right)$
-$\bm{Y}=\left(\bm{C}\bm{B}^{T}\right)\bm{X}\quad\Rightarrow\quad\bm{Y}=\bm{C}\left(\bm{B}^{T}\bm{X}\right)$
-To integrate causal masking without compromising linear computational complexity, the block properties of the 1-SS matrix $\bm{L}$ can be leveraged in conjunction with segment accumulation.
-This approach facilitates the implementation of causal linear attention, referred to as Structured Masked Attention (SMA) (SSD) in SSD.
+(8) $\bm{Y}=\bm{L}\circ\bm{C}\bm{B}^{\top}\bm{X}=\bm{L}\circ\bm{Q}\bm{K}^{\top}\bm{V}.$
+This concept is referred to as Duality in the SSD framework (SSD2024).
+By referring to the tensor contraction order in the calculation of linear attention (LinearAttention2020)—specifically using the associative property of matrix multiplication where $K^{T}V$ is computed first and subsequently left-multiplied by $Q$—the computational complexity is reduced from quadratic to linear in the sequence length without considering causal masks, a property that SSD also achieves.
+$\bm{Y}=\left(\bm{Q}\bm{K}^{T}\right)\bm{V}\quad\Rightarrow\quad\bm{Y}=\bm{Q}\left(\bm{K}^{T}\bm{V}\right),$
+$\bm{Y}=\left(\bm{C}\bm{B}^{T}\right)\bm{X}\quad\Rightarrow\quad\bm{Y}=\bm{C}\left(\bm{B}^{T}\bm{X}\right),$
+To integrate causal masked matrix without compromising linear computational complexity, the block properties of the 1-SS matrix $\bm{L}$ can be leveraged in conjunction with segment accumulation.
+This approach facilitates the implementation of causal linear attention, referred to as Structured Masked Attention (SMA) (SSD2024) in SSD.
 Since the SSD architecture is derived from the Mamba model, models based on the SSD architecture are referred to as Mamba2.
@@
 3. METHOD
+In this section, we offer a detailed introduction to our proposed framework, TiM4Rec.
+Following an overview of the overall structure of TiM4Rec, we delve into the reasons behind the performance degradation of SSD compared to SSM in low-dimensional scenarios within sequential recommendation tasks.
+Subsequently, we explain how our time-aware enhancement method, utilizing our proposed Time-aware Masked Matrix with linear computational complexity, compensates for this performance degradation.
+3.1. Problem Formulation and Method Overview
 Figure 2.
@@
 The overall framework of TiM4Rec, which is based on time-aware enhancement, enhances the performance of the SSD architecture in the domain of sequential recommendation by effectively processing the time differences in interactions between items.
-In this section, we offer a detailed introduction to our proposed framework, TiM4Rec.
-Following an overview of the overall structure of TiM4Rec, we delve into the reasons behind the performance degradation of SSD compared to SSM in low-dimensional scenarios within sequential recommendation tasks.
-Subsequently, we explain how our time-aware enhancement method, utilizing our proposed time-aware masked matrix with linear computational complexity, compensates for this loss.
-3.1. Overview of TiM4Rec
-As shown in Figure 2, the overall logic of TiM4Rec adheres to the description in section 2.1, However, an additional user interaction unix timestamp sequence $\mathcal{T}^{u_{k}}=\left\{t_{1},t_{2},\cdots,t_{L}\right\}\in\mathbb{R}^{L}$ is incorporated into the input data, with $\mathcal{T}^{u_{k}}$ corresponding one to one with the user interation sequence $\mathcal{S}^{u_{k}}$, where $L$ representing interaction sequence length.
-The sequence $\mathcal{S}^{u_{k}}$ can be transformed into an input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ composed of dense vectors by the embedding table $\mathcal{E}$.
-Subsequently, a displacement subtraction is performed on sequence $\mathcal{T}^{u_{k}}$, and a zero is appended to the begining to generate the intearaction time difference sequence $\mathcal{D}^{u_{k}}=\left\{d_{1},d_{2},\cdots,d_{L}\right\}\in\mathbb{R}^{L}$, which be used in later stages.
-The Time-aware SSD Block enhances the model’s ability to capture potential user interests by integrating sequence $\mathcal{D}^{u_{k}}$ into a discretized SSD kernel.
-For the sake of convenience in the subsequent discussion, the superscript $u_{k}$ be omitted.
-3.2. Time Difference Processing
-As mentioned above, for the input interaction timestamp sequence $\mathcal{T}$, we transform it into an interaction time difference sequence $\mathcal{D}$ using the following method:
-(5) $\mathcal{D}_{i}=\begin{cases}0,\quad&i=1\\ \mathcal{T}_{i}-\mathcal{T}_{i-1},\quad&i\in\left[2,L\right]\end{cases}$
-Subsequently, we apply dropout (srivastava2014dropout) and layer normalization (ba2016layer) to sequence $\mathcal{D}$ in order to eliminate data noise and optimize the data distribution for subsequent analysis.
-3.3. Analyzing the Masked Matrix in SSD
+Let $\mathcal{U}=\left\{u_{1},u_{2},\cdots,u_{|\mathcal{U}|}\right\}$ be the set of users and $\mathcal{V}=\left\{v_{1},v_{2},\cdots,v_{|\mathcal{V}|}\right\}$ be the set of items.
+For any user $u_{k}$, there exists an interaction sequence $\mathcal{S}^{u_{k}}=\left\{v_{1}^{u_{k}},v_{2}^{u_{k}},\cdots,v_{T}^{u_{k}}\right\}\in\mathbb{R}^{T}$ order by the user interaction unix timestamp sequence $\mathcal{T}^{u_{k}}=\left\{t_{1},t_{2},\cdots,t_{T}\right\}\in\mathbb{R}^{T}$, where $v_{i}^{u_{k}}$ represents the i-th item of user $u_{k}$ interaction and $T$ representing interaction sequence length.
+Sequential recommendation methodologies leverage the user’s historical interaction sequence $\mathcal{S}^{u_{k}}$ to forecast the next potential interaction item $v_{T+1}$ for the user $u_{k}$.
+Without loss of generality, we extract the preference representation $p^{u_{k}}$ for user $u_{k}$ from the historical interaction sequence $\mathcal{S}^{u_{k}}$ and the timestamp sequence $\mathcal{T}^{u_{k}}$, thereby predicting the probability distribution $\hat{\mathcal{Y}}\left(v_{T+1}|p^{u_{k}}\right)$ for the next potential interaction item.
+For convenience in the subsequent discussion, the superscript $u_{k}$ be omitted.
+As illustrated in Fig.
+2, the proposed TiM4Rec is a time-aware sequential recommendation model that leverages the SSD architecture through the Time-aware SSD blocks.
+The TiM4Rec model consists of multiple layers of Time-aware SSD Layers, each composed of a Time-aware SSD Block and a Feed Forward Network (FFN) (Transformer2017).
+We incorporate a special time difference processing method that allows the SSD kernel to integrate time difference information.
+3.2. Data Preprocessing
+3.2.1. Embedding Layer
+In a manner analogous to previous models, we establish a learnable embedding table $\mathbb{E}=\left\{E_{1},E_{2},\cdots,E_{|\mathcal{V}|}\right\}\in\mathbb{R}^{|\mathcal{V}|\times D}$, the item IDs are converted from a sparse vector representation to a dense vector representation, where $E_{j}$ denotes the potential dense vector representation of the item $v_{j}$.
+The interaction sequence $\mathcal{S}$ can be transformed into an input sequence $\bm{X}\in\mathbb{R}^{T\times D}$ composed of dense vectors by the embedding table $\mathbb{E}$.
+Subsequently, we enhance the embedding table $\mathbb{E}$ representation learning process by employing dropout (Dropout2014) and layer normalization (LayerNorm2016) techniques.
+These methods mitigate overfitting and stabilize the training dynamics, thereby improving the robustness and generalization capabilities of the learned embeddings.
+3.2.2. Time Difference Processing
+To integrate the interactive timestamp sequence $\mathcal{T}$ into the SSD computation kernel and enhance the model’s ability to capture latent user interests, we perform differential processing on $\mathcal{T}$.
+The specific mechanisms of this integration will be discussed shortly.
+The interactive timestamp sequence $\mathcal{T}$ undergoes a displacement subtraction, with an initial zero appended, resulting in the time difference sequence $\mathcal{D}=\left\{d_{0},d_{1},\cdots,d_{T-1}\right\}\in\mathbb{R}^{T}$.
+The computation process is as follows:
+(9) $\mathcal{D}_{i}=\begin{cases}0,\quad&i=0\\ \mathcal{T}_{i+1}-\mathcal{T}_{i},\quad&i\in\left[1,T-1\right]\end{cases}.$
+Subsequently, we apply dropout (Dropout2014) and layer normalization (LayerNorm2016) to sequence $\mathcal{D}$ in order to eliminate data noise and optimize the data distribution for subsequent analysis.
+3.3. Time-Aware Linear Complexity Enhancement
+3.3.1. Analyzing the Masked Matrix in SSD
 As discussed in section 2.4, the critical advancement from SSM to SSD resides in the scalarization of the transition matrix $\bm{A}$.
-When considering this modification from an attention perspective, the central aspect involves the alterations in the mask matric $\bm{L}$.
+When considering this modification from an attention perspective, the central aspect involves the alterations in the masked matrix $\bm{L}$.
 Although the recursive operational constraints of matrix $\bm{A}$ in SSM prevent the direct and precise formulation of the masked matrix within Eq.
-3, it remains feasible to examine an equivalent expression for the final masked matrix:
-$\bm{L}_{\text{SSM}}=\begin{bmatrix}I\\ \bm{A}_{1}&I\\ \bm{A}_{2}\bm{A}_{1}&\bm{A}_{2}&I\\ \vdots&\vdots&\ddots&\ddots\\ \bm{A}_{1:i}^{\times}&\bm{A}_{2:i}^{\times}&\cdots&\bm{A}_{i-1:i}^{\times}&I\end{bmatrix}$
-As illustrated in figure 3, in contrast to the masked matrix of SSD presented in eq.
-4, the fundamental change in the masked matrix $L_{\text{SSM}}$ is that its mask coefficients are applied in a point-to-point manner on the individual item $\bm{E}_{i}=\left[e_{1},e_{2},\cdots,e_{n}\right]\in\mathbb{R}^{N}$ (The $\bm{A}$ matrix follows its definitions in SSM and SSD, respectively.):
+5, it remains feasible to examine an equivalent expression for the final masked matrix:
+(10) $\bm{L}_{\text{SSM}}=\begin{bmatrix}I\\ \bm{A}_{1}&I\\ \bm{A}_{2}\bm{A}_{1}&\bm{A}_{2}&I\\ \vdots&\vdots&\ddots&\ddots\\ \bm{A}_{1:i}^{\times}&\bm{A}_{2:i}^{\times}&\cdots&\bm{A}_{i-1:i}^{\times}&I\end{bmatrix},$
+As illustrated in Fig.
+3, in contrast to the masked matrix of SSD presented in Eq.
+6, the fundamental change in the masked matrix $L_{\text{SSM}}$ is that its mask coefficients are applied in a point-to-point manner on the individual item $\bm{E}_{i}=\left[e_{1},e_{2},\cdots,e_{n}\right]\in\mathbb{R}^{D}$ (The $\bm{A}$ matrix follows its definitions in SSM and SSD, respectively.):
 […]
@@
 $\bm{L}_{MSA}=\begin{bmatrix}1\\ 1&1\\ 1&1&1\\ \vdots&\vdots&\ddots&\ddots\\ 1&1&\cdots&1&1\end{bmatrix},$
-MSA stands for Vanilla Mask Self Attention Mechanism
-However, owing to the semi-separable matrix structure inherent in the masked matrix of SSD, there is a pronounced attenuation effect on the interaction items that appear earlier in the sequence.
+MSA stands for Vanilla Masked Self Attention Mechanism
+However, owing to the 1-SS matrix structure inherent in the masked matrix of SSD, there is a pronounced attenuation effect on the interaction items that appear earlier in the sequence.
 This feature aligns SSD more closely with the actual user interaction patterns observed in real-world recommendation systems.
@@
 For an individual item, SSM can be viewed as an equivalent point-to-point mask, whereas SSD resembles the global mask utilized in Transformer architectures.
-Observing the differences between the equivalent mask matrices of SSD and SSM, it can be seen that the core of their mask matrices is recursively generated based on the state transition matrix $\bm{A}$.
+Observing the differences between the equivalent masked matrices of SSD and SSM, it can be seen that the core of their mask matrices is recursively generated based on the state transition matrix $\bm{A}$.
 As described in section 2.4, the only difference is that SSD achieves more efficient matrix multiplication calculations by scalar quantization of matrix $\bm{A}$.
 Therefore, addressing the loss of the masked matrix caused by the scalar quantization of matrix $\bm{A}$ while maintaining the efficient matrix multiplication calculation of SSD has become the key to improvement.
-3.4. Time-aware Masked Matrix
-Let’s reconsider how the time difference matrix should be expressed in the context of attention mechanisms.
-The method proposed by TiSASRec (TiSASRec) is centered around acquiring the time position encoding for each item in the interaction sequence through interaction timestamps, ultimately leading to the generation of the corresponding attention bias matrix based on these encodings.
-Given that the Mamba architecture does not depend on positional encoding—as substantiated by the findings in Mamba4Rec (mamba4rec) and SSD4Rec (ssd4rec)—it is essential to directly explore the equivalent representation of the TiSASRec approach.
-Specifically, we should investigate how the resulting attention bias matrix encapsulates the temporal differences in product interactions.
-For a sequence of interactive timestamps $\mathcal{T}$, we can derive the relative time difference matrix $\hat{\mathcal{D}}$ between each product through the following transformation:
-(6) $\hat{\mathcal{D}}=\begin{bmatrix}\hat{d}_{11}\\ \mathbf{\underline{\hat{d}_{21}}}&\hat{d}_{22}\\ \hat{d}_{31}&\mathbf{\underline{\hat{d}_{32}}}&\hat{d}_{33}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{d}_{L1}&\hat{d}_{L2}&\cdots&\mathbf{\underline{\hat{d}_{LL-1}}}&\hat{d}_{LL}\end{bmatrix},\quad\hat{d}_{mn}=t_{m}-t_{n}$
+3.3.2. Time-aware Structured Masked Matrix
+Let’s reconsider the time-aware enhancement approach of TiSASRec (TiSASRec2020).
+The essence of this method lies in embedding the time difference matrix and processing it through the attention mechanism to generate a time-aware attention score matrix, which is then integrated into the final attention score matrix.
+If we regard the generated time-aware attention score matrix as a special type of masked matrix, it aligns with our previous analysis: employing a more refined mask matrix to extract potential user interest features can lead to improved recommendation performance.
+However, in the TiSASRec method, the computational steps involved in generating the time-aware masked matrix by embedding the time difference matrix and processing it through the attention mechanism incur substantial computational overhead.
+Therefore, it is imperative to identify a direct and efficient approach to convert the time difference matrix into a time-aware masked matrix.
+For a sequence of interactive timestamps $\mathcal{T}$ (The definition is the same as in section 3.1), we can derive the relative time difference matrix $\hat{\mathcal{D}}$ between each product through the following transformation:
+(11) $\hat{\mathcal{D}}=\begin{bmatrix}\hat{d}_{11}\\ \mathbf{\underline{\hat{d}_{21}}}&\hat{d}_{22}\\ \hat{d}_{31}&\mathbf{\underline{\hat{d}_{32}}}&\hat{d}_{33}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{d}_{T1}&\hat{d}_{T2}&\cdots&\mathbf{\underline{\hat{d}_{TT-1}}}&\hat{d}_{TT}\end{bmatrix},\quad\hat{d}_{mn}=t_{m}-t_{n},$
 Directly applying matrix $\hat{\mathcal{D}}$ to the masked matrix $\bm{L}$ (as described in Eq.
-4) compromises the semi-separable matrix property of $\bm{L}$.
+6) compromises the 1-SS matrix property of $\bm{L}$.
 This compromise prevents us from leveraging the efficient matrix computation techniques offered by the SSD architecture, consequently resulting in a return to quadratic computational complexity.
@@
 When we construct a new matrix $\mathcal{D}$ by selecting only the values of the underlined elements in matrix $\hat{\mathcal{D}}$, we effectively adopt the form described in Eq.
-5.
+9.
 This approach allows us to integrate matrix $\mathcal{D}$ into the masked matrix $\bm{L}$ with linear complexity.
 Initially, we apply matrix $\mathcal{D}$ to the scalarization matrix $\bm{A}$, which results in the corresponding positions in the masked matrix $\bm{L}$ containing the relevant time difference information (Due to the presence of layer normalization, the actual value of $d_{1}$ will be non-zero.):
-$\hat{\bm{A}}_{i}=\hat{a}_{i}=\mathcal{D}_{i}\cdot\bm{A}_{i}=\hat{d}_{ii-1}\cdot a_{i}=d_{i}\cdot a_{i}$
-$\hat{\bm{A}}_{m:n}\coloneqq\hat{a}_{n}\hat{a}_{n-1}\cdots\hat{a}_{m}=a_{n}d_{n}\cdot a_{n-1}d_{n-1}\cdots a_{m}d_{m}$
-(7) $\bm{L}=\begin{bmatrix}1\\ \hat{a}_{1}&1\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&1\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&1\end{bmatrix}$
-This can be regarded as an equivalent approximation of $d_{ij}$ as follows:
-$\hat{d}_{ij}\approx\hat{d}_{ii-1}\hat{d}_{i-1i-2}\cdots\hat{d}_{j+1j}=d_{i}d_{i-1}\cdots d_{j}$
-In the context of SSD kernel computations, multiplication is effectively converted into accumulation operations via the logarithm function.
-This transformation offers advantageous support for the aforementioned approximation of $\hat{d}_{ij}$.
-$\log\left(d_{i}d_{i-1}\cdots d_{j}\right)=\log\left(d_{i}\right)+\log\left(d_{i-1}\right)+\cdots+\log\left(d_{j}\right)$
-Through the approximation method mentioned above, we accomplish the time-aware enhancement of the SSD architecture.
-This is achieved without disrupting the semi-separable property of the masked matrix $\bm{L}$.
+$\hat{\bm{A}}_{i}=\hat{a}_{i}=\mathcal{D}_{i}\cdot\bm{A}_{i}=\hat{d}_{ii-1}\cdot a_{i}=d_{i}\cdot a_{i},$
+$\hat{\bm{A}}_{m:n}\coloneqq\hat{a}_{n}\hat{a}_{n-1}\cdots\hat{a}_{m}=a_{n}d_{n}\cdot a_{n-1}d_{n-1}\cdots a_{m}d_{m},$
+(12) $\bm{L}=\begin{bmatrix}\hat{a}_{0}\\ \hat{a}_{1}&\hat{a}_{0}\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&\hat{a}_{0}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&\hat{a}_{0}\end{bmatrix},$
+We refer to $\bm{L}$ as a Time-aware Structured Masked Matrix.
+This can be regarded as an equivalent approximation of $\hat{d}_{ij}$ as follows (Notably, the computation in Eq.
+12 employs the segment accumulation, consistent with the 1-SS matrix computation in SSD (SSD2024), to maintain efficient computational performance.):
+$\hat{d}_{ij}\approx\hat{d}_{ii-1}\hat{d}_{i-1i-2}\cdots\hat{d}_{j+1j}=d_{i}d_{i-1}\cdots d_{j},$
+Through the approximation method mentioned above, we accomplish the time-aware enhancement of the SSD architecture. This is achieved without disrupting the 1-SS property of the masked matrix $\bm{L}$.
+Figure 4.
+In comparing our proposed method with the time-aware enhancement approach based on the TiSASRec paradigm, we emphasize the distinctions in how time difference information is transformed into attention score matrices.
+The time difference matrix is defined in the upper left corner, as outlined in Eq.
+11.
+By directly integrating the time difference matrix into the masked matrix using the aforementioned method, computational efficiency is maintained.
+As illustrated in Figure.
+4, unlike the time-aware augmentation methods based on the TiSASRec paradigm (TiSASRec2020), our approach circumvents the need to elevate time difference information to high-dimensional processing.
+Instead, it employs a straightforward scalar-level sequence transformation (which will be detailed shortly) to effectively capture personalized user interaction patterns.
+This enhanced time difference information is then utilized to construct a Time-aware Structured Masked Matrix $\bm{L}$ that adheres to the property of 1-SS matrix, integrating seamlessly into the SSD kernel.
+This approach addresses the dual challenges of computational efficiency and integration into the SSD kernel.
 3.4. Time-aware SSD Block
+In the following section, we will elaborate on the integration of the Time-aware Structured Masked Matrix into the SSD kernel.
 Despite providing a detailed derivation of the SSD in Section 2.4, the actual code implementation of the SSD kernel retains the discretization process.
@@
 Specifically, for matrix $\overline{\bm{A}}$, the implementation approximates the exponential operation by directly computing $\overline{\bm{A}}\approx\Delta\bm{A}$.
-Meanwhile, for matrix $\overline{\bm{B}}$, given the Maclaurin series expansion of $e^{x}$ as $1+x+\frac{x^{2}}{2}+\cdots$, the following approximation can be made:
-$\overline{\bm{B}}$ $\approx\left(\Delta\bm{A}\right)^{-1}\left(\bm{I}+\Delta\bm{A}-\bm{I}\right)\cdot\Delta\bm{B}$
+Meanwhile, for matrix $\overline{\bm{B}}$, given the Maclaurin series expansion of $\exp(x)$ as $1+x+\frac{x^{2}}{2}+\cdots$, the following approximation can be made:
+(13) $\overline{\bm{B}}$ $\approx\left(\Delta\bm{A}\right)^{-1}\left(\bm{I}+\Delta\bm{A}-\bm{I}\right)\cdot\Delta\bm{B}$
 $=\left(\Delta\bm{A}\right)^{-1}\Delta\bm{A}\cdot\Delta\bm{B}$
-$=\Delta\bm{B}$
-In this approximation, we establish the following definitions: the state transition coefficient $\bm{A}\in\mathbb{R}^{1}$.
+$=\Delta\bm{B},$
+Under this approximation, the original state transition matrix $\bm{A}\in\mathbb{R}^{T}$ in the SSD can be transformed into a scalar coefficient $A\in\mathbb{R}^{1}$, which when multiplied by the discretization parameter $\Delta\in\mathbb{R}^{T}$, yields the discretized state transition matrix $\overline{\bm{A}}\in\mathbb{R}^{T}$.
+This approximation process optimizes computational efficiency while maintaining maximal accuracy.
+Given that this approach is empirically established by Gu et al.
+(SSD2024), we will not delve into further details here.
+Building upon the aforementioned consensus, we now proceed to a comprehensive derivation of the Time-aware SSD Block method.
 By inputting $\bm{X}\in\mathbb{R}^{T\times D}$, we derive matrices $\bm{B}\in\mathbb{R}^{T\times N}$ and $\bm{C}\in\mathbb{R}^{T\times N}$, along with the discretization parameter $\Delta\in\mathbb{R}^{T}$:
-$\bm{X},\bm{B},\bm{C},\Delta=\bm{X}W+b$
+(14) $\bm{X},\bm{B},\bm{C},\Delta=\bm{X}W+b,$
 Where the weight matrix $W\in\mathbb{R}^{D\times(D+2N+1)}$ and the bias matrix $b\in\mathbb{R}^{D+2N+1}$.
 Subsequently, a causal convolution transformation is applied to the matrices $\bm{X}$, $\bm{B}$ and $\bm{C}$:
-$\bm{X}_{t},\bm{B}_{t},\bm{C}_{t}$ $=\sigma\left[(\bm{X}_{t},\bm{B}_{t},\bm{C}_{t})^{\top}*\omega\right]$
-$\text{where}\quad\mathcal{Q}_{t}=\mathcal{P}_{t}*\omega$ $\coloneqq\sum_{m=0}^{K-1}\mathcal{P}_{\max\left(t-m,0\right)}\cdot\omega_{m}$
+(15) $\bm{X}_{t},\bm{B}_{t},\bm{C}_{t}=\sigma\left[(\bm{X}_{t},\bm{B}_{t},\bm{C}_{t})^{\top}*\omega\right],$
+$\text{where}\quad\mathcal{Q}_{t}=\mathcal{P}_{t}*\omega\coloneqq\sum_{m=0}^{K-1}\mathcal{P}_{\max\left(t-m,0\right)}\cdot\omega_{m},$
 Where the $\omega\in\mathbb{R}^{K}$ is convolution kernel and the $\sigma$ represents non-linear activation functions.
@@
 Therefore, we initially apply a time-varying transformation to the time difference matrix using the following transformation:
-$\alpha^{\mathcal{D}}=\sigma\left[\sigma\left(\mathcal{D}W_{1}^{\mathcal{D}}+b_{1}^{\mathcal{D}}\right)W_{2}^{\mathcal{D}}+b_{2}^{\mathcal{D}}\right]$
-$\mathcal{D}=\alpha^{\mathcal{D}}\cdot\mathcal{D}$
+(16) $\alpha^{\mathcal{D}}$ $=\sigma\left[\sigma\left(\mathcal{D}W_{1}^{\mathcal{D}}+b_{1}^{\mathcal{D}}\right)W_{2}^{\mathcal{D}}+b_{2}^{\mathcal{D}}\right],$
+$\mathcal{D}$ $=\alpha^{\mathcal{D}}\cdot\mathcal{D},$
 Where the weight matrix $W_{1}^{\mathcal{D}},W_{2}^{\mathcal{D}}\in\mathbb{R}^{T\times T}$ and the bias matrix $b_{1}^{\mathcal{D}},b_{2}^{\mathcal{D}}\in\mathbb{R}^{T}$.
 Considering the convergence of interaction time differences within a short time frame, we apply causal convolution to the time difference matrix $\mathcal{D}$ to enhance the implicit features of user interest points.
-$\mathcal{D}_{t}=\sigma\left(\mathcal{D}_{t}*\omega^{\mathcal{D}}\right)\coloneqq\sigma\left(\sum_{m=0}^{K-1}\mathcal{D}_{\max\left(t-m,0\right)}\cdot\omega^{\mathcal{D}}_{m}\right)$
+(17) $\mathcal{D}_{t}=\sigma\left(\mathcal{D}_{t}*\omega^{\mathcal{D}}\right)\coloneqq\sigma\left(\sum_{m=0}^{K-1}\mathcal{D}_{\max\left(t-m,0\right)}\cdot\omega^{\mathcal{D}}_{m}\right),$
 Next, we integrate the enhanced time difference matrix $\mathcal{D}$ with the discrete parameter $\Delta$, and obtain matrices $\overline{\bm{A}}$ and $\overline{\bm{B}}$ through the following transformation:
-$\hat{\Delta}=\Delta\cdot\mathcal{D},\quad\overline{\bm{A}}=\hat{\bm{A}}=\hat{\Delta}\cdot\bm{A},\quad\overline{\bm{B}}=\hat{\Delta}\cdot\bm{B}$
-Finally, the following equation can be derived to map the input $\bm{X}$ to the output $\bm{Y}\in\mathbb{R}^{T\times D}$($\bm{L}$ refers to eq. 7.):
-$\bm{Y}=\bm{L}\circ\bm{C}\overline{\bm{B}}^{\top}\bm{X}=\bm{L}\circ\bm{C}\bm{B}^{\top}\left(\hat{\Delta}^{\top}\bm{X}\right)$
-Due to the preservation of the semi-separable matrix property of $\bm{L}$, the matrix computation acceleration method discussed in Section 2.3 remains applicable.
-To extract deeper user interest features, multiple layers of Time-aware SSD blocks are stacked.
-Consequently, after the Time-aware SSD, the features are transformed by adding a Feed Forward Network layer to adapt to the next semantic space.
-It is noteworthy that to adapt the time difference vector $\mathcal{D}$ to the feature semantic space of the next layer, gate residual (he2016deep) processing is applied to the input time difference vector $\mathcal{D}$ for the subsequent layer.
+(18) $\hat{\Delta}=\Delta\cdot\mathcal{D},\quad\overline{\bm{A}}=A\cdot\hat{\Delta},\quad\overline{\bm{B}}=\hat{\Delta}\cdot\bm{B},$
+Following the methodology outlined in Section 3.3.2, we employ the following transformations to construct the Time-aware Structured Masked Matrix $\bm{L}$:
+$\hat{\bm{A}}_{i}=\hat{a}_{i}=\overline{\bm{A}}_{i}=A\cdot\hat{\Delta}_{i}=A\cdot\Delta_{i}\cdot d_{i},$
+(19) $\bm{L}=\begin{bmatrix}\hat{a}_{0}\\ \hat{a}_{1}&\hat{a}_{0}\\ \hat{a}_{2}\hat{a}_{1}&\hat{a}_{2}&\hat{a}_{0}\\ \vdots&\vdots&\ddots&\ddots\\ \hat{a}_{t-1}\ldots\hat{a}_{1}&\hat{a}_{t-1}\ldots\hat{a}_{2}&\cdots&\hat{a}_{t-1}&\hat{a}_{0}\end{bmatrix},$
+Finally, the following equation can be derived to map the input sequence $\bm{X}$ and $\mathcal{D}$ to the output $\bm{Y}\in\mathbb{R}^{T\times D}$:
+(20) $\bm{Y}=TiSSD\left(\bm{X},\mathcal{D}\right)\coloneqq\bm{L}\circ\bm{C}\overline{\bm{B}}^{\top}\bm{X}=\bm{L}\circ\bm{C}\bm{B}^{\top}\left(\hat{\Delta}^{\top}\bm{X}\right),$
+Due to the preservation of the 1-SS matrix property of $\bm{L}$, the matrix computation acceleration method discussed in Section 2.4 remains applicable.
+Multiple layers of Time-aware SSD blocks are stacked to extract deeper user interest features.
+Consequently, after the Time-aware SSD, the features are transformed by adding an FFN (Transformer2017) layer to adapt to the next semantic space.
+It is noteworthy that to adapt the time difference vector $\mathcal{D}$ to the feature semantic space of the next layer, gate residual (ResNet2016) processing is applied to the input time difference vector $\mathcal{D}$ for the subsequent layer.
 3.5. Prediction Layer
-By utilizing the Time-aware SSD Block, input $\bm{X}$ can be transformed into a user interest feature sequence $\bm{P}\in\mathbb{R}^{L\times D}$.
-The last element $p_{t}$ of the feature sequence $\bm{P}$ is then extracted as the interest feature corresponding to the current user interaction sequence.
-$\hat{\mathcal{Y}}=Softmax\left(p_{t}\cdot\mathcal{E}\right)$
-As stated in Section 2.1, the predicted scores $\hat{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{V}|}$ for all items are obtained by calculating the inner product of $p_{t}$ and the embedding table of all items.
-However, a notable difference is that the actual score calculation incorporates the Softmax function, even though the Softmax operation does not influence the ranking of the scores.
+By utilizing the Time-aware SSD Layer, input $\bm{X}$ can be transformed into a user interest feature sequence $\bm{P}\in\mathbb{R}^{T\times D}$.
+The last element $p\in\mathbb{R}^{D}$ of the feature sequence $\bm{P}$ is then extracted as the interest feature corresponding to the current user interaction sequence.
+Afterward, we compute the dot product between the user interest feature $p$ and the embeddings of all items $\mathbb{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$.
+This result is then scaled using the Softmax function to obtain the interaction probability scores $\hat{\mathcal{Y}}\left(v_{T+1}|p\right)\in\mathbb{R}^{|\mathcal{V}|}$ for each item from the user.
+(21) $\hat{\mathcal{Y}}\left(v_{T+1}|p\right)=Softmax\left(p\cdot\mathbb{E}^{T}\right),$
+In alignment with previous work (SASRec2018), we optimize the model parameters using the Cross Entropy loss function.
+Let us denote the true interaction distribution for user $u_{k}$ as $g^{u_{k}}\in\mathbb{R}^{|\mathcal{V}|}$; then, the following loss function is defined:
+(22) $\mathcal{L}=-\sum_{u_{k}}^{|\mathcal{U}|}\sum_{i}^{|\mathcal{V}|}g^{u_{k}}_{i}\ln\left(\hat{\mathcal{Y}}^{u_{k}}_{i}\right)$
 3.6. Complexity Analysis
-Models based on the Vanilla Transformer, such as SASRec, require a total of $O(L^{2}N)$ floating-point operations (FLOPs).
-In contrast, thanks to the computational improvements provided by Mamba, models based on SSM and SSD, such as Mamba4Rec and SSD4Rec, only require $O(LN^{2})$ FLOPs.
-In practical application scenarios of sequential recommendation, including news and product recommendations, it is often the case that $L\gg N$.
+Models based on the Vanilla Transformer (Transformer2017), such as SASRec (SASRec2018), require a total of $O(L^{2}N)$ floating-point operations (FLOPs).
+Models based on Linear Recurrent Units (LRU2023), such as LRURec (LRURec2024), require $O(\log(T)N^{2})$ FLOPs.
+In contrast, thanks to the computational improvements provided by Mamba, models based on SSM (Mamba2023) and SSD (SSD2024), such as Mamba4Rec (Mamba4Rec2024) and SSD4Rec (SSD4Rec2024), only require $O(TN^{2})$ FLOPs.
+Notably, although the computational complexity of the SSM and SSD architectures is identical, the SSD architecture can efficiently leverage matrix multiplication for computations, resulting in faster actual training times compared to the SSM architecture.
+In practical application scenarios of sequential recommendation, including news and product recommendations, it is often the case that $T\gg N$.
 Thus, sequential recommendation models based on Mamba demonstrate superior advantages.
-TiSASRec enhances its performance by integrating time-aware capabilities via time difference encoding.
-Despite the theoretical FLOPS remaining in the order of $O(L^{2}N)$, the practical computational and memory demands exceed those of SASRec by more than four times.
+TiSASRec (TiSASRec2020) enhances its performance by integrating time-aware capabilities via time difference encoding.
+Despite the theoretical FLOPS remaining in the order of $O(T^{2}N)$, the practical computational and memory demands exceed those of SASRec by more than four times.
 This substantial increase is attributed to the inherent computational constraints associated with time difference encoding.
 Consequently, while TiSASRec demonstrates improved time sensitivity, it necessitates a significantly higher computational overhead and memory allocation compared to SASRec, posing potential challenges in scalability and efficiency.
-In contrast, TiM4Rec preserves the benefits of the SSD architecture by employing a time difference masked matrix characterized by linear computational complexity.
-This approach maintains the computational efficiency, demanding only $O(LN^{2})$ FLOPs, and incurs only a modest increase in additional computational overhead.
+In contrast, TiM4Rec preserves the benefits of the SSD architecture by employing a Time-aware Structured Masked Matrix characterized by linear computational complexity.
+This method preserves computational efficiency, requiring only $O(TN^{2})$FLOPs, and demands only an additional $O(T)$ level of FLOPs.
 Notably, the generation of the time-aware masked matrix in TiM4Rec is solely based on scalar computations.
 This characteristic ensures that the additional computational complexity does not escalate in high-dimensional modeling scenarios, thereby preserving the superior matrix operation efficiency inherent to the SSD4Rec architecture.
-Table 1. Comparison of computational complexity between representative sequential recommendation methods, where ”M.M.” refers to ”Matrix Multiplication”.
-Architecture FLOPs Memory M.M.
-SASRec Attention $L^{2}N$ $L^{2}N$
-Mamba4Rec SSM $LN^{2}$ $LN^{2}$
-SSD4Rec SSD $LN^{2}$ $LN$
-TiSASRec Attention $L^{2}N$ $L^{2}N$
-TiM4Rec SSD $LN^{2}$ $LN$
+Our summarized results are shown in Table 1.
+Table 1. Comparison of computational complexity between representative sequential recommendation methods, where ”M.M.” refers to ”Matrix Multiplication”, ”T.F” represents ”Time-aware additional FLOPs”
+Architecture FLOPs T.F M.M.
+SASRec Attention $T^{2}N$
+LRURec LRU $\log(T)N^{2}$
+Mamba4Rec SSM $TN^{2}$
+SSD4Rec SSD $TN^{2}$
+TiSASRec Attention $T^{2}N$ $T^{2}N$
+TiM4Rec SSD $TN^{2}$ $T$
 4. Experiments
-Table 2.
-Sequential recommendation performance on the ML-1M, Beauty, and KuaiRand datasets.
-Here, R@k refers to Recall@K, N@K denotes NDCG@K, and M@K represents MRR@K.
-The best and second-best results are indicated in bold and underlined, respectively.
-”Improv.” signifies the performance improvement of TiM4Rec compared to the second-best results.
-The asterisk in SSD4Rec* denotes that this is our own replicated version.
-Model Type & Model $\rightarrow$ CNN RNN Transformer SSM SSD Improv.
-Dataset Metric Caser GRU4Rec SASRec BERT4Rec TiSASRec Mamba4Rec SSD4Rec* TiM4Rec
-MovieLens-1M R@10 0.2156 0.2985 0.3060 0.2800 0.3147 0.3253 0.3199 0.3310 1.75%
-R@20 0.3083 0.3937 0.4050 0.3853 0.4250 0.4354 0.4240 0.4338 -0.30%
-R@50 0.4515 0.5397 0.5455 0.5377 0.5608 0.5707 0.5649 0.5770 1.10%
-N@10 0.1157 0.1700 0.1754 0.1584 0.1834 0.1891 0.1841 0.1932 2.17%
-N@20 0.1391 0.1941 0.2005 0.1852 0.2106 0.2169 0.2104 0.2194 1.15%
-N@50 0.1675 0.2232 0.2285 0.2154 0.2376 0.2436 0.2384 0.2477 1.68%
-M@10 0.0853 0.1307 0.1357 0.1214 0.1424 0.1474 0.1425 0.1512 2.58%
-M@20 0.0917 0.1373 0.1425 0.1288 0.1498 0.1551 0.1498 0.1584 2.13%
-M@50 0.0963 0.1420 0.1471 0.1337 0.1542 0.1593 0.1542 0.1629 2.26%
-Amazon Beauty R@10 0.0402 0.0563 0.0851 0.0352 0.0802 0.0838 0.0806 0.0854 0.35%
-R@20 0.0618 0.0856 0.1194 0.0563 0.1147 0.1185 0.1146 0.1204 0.84%
-R@50 0.1090 0.1376 0.1759 0.0999 0.1734 0.1802 0.1710 0.1800 -0.11%
-N@10 0.0199 0.0297 0.0425 0.0172 0.0405 0.0435 0.0423 0.0446 2.53%
-N@20 0.0253 0.0371 0.0511 0.0225 0.0491 0.0522 0.0509 0.0533 2.11%
-N@50 0.0346 0.0473 0.0623 0.0311 0.0605 0.0644 0.0620 0.0651 1.09%
-M@10 0.0138 0.0217 0.0294 0.0117 0.0282 0.0311 0.0306 0.0321 3.22%
-M@20 0.0153 0.0237 0.0317 0.0132 0.0306 0.0335 0.0329 0.0345 2.99%
-M@50 0.0167 0.0253 0.0335 0.0145 0.0325 0.0354 0.0347 0.0363 2.54%
-Kuai Rand R@10 0.0801 0.1020 0.1055 0.0938 0.1057 0.1094 0.1055 0.1109 1.37%
-R@20 0.1344 0.1659 0.1704 0.1537 0.1710 0.1768 0.1717 0.1774 0.34%
-R@50 0.2561 0.3017 0.3074 0.2873 0.3060 0.3154 0.3088 0.3202 1.52%
-N@10 0.0395 0.0564 0.0584 0.0510 0.0590 0.0608 0.5880 0.0611 0.49%
-N@20 0.0531 0.0724 0.0747 0.0660 0.0753 0.0777 0.0754 0.0779 0.26%
-N@50 0.0770 0.0911 0.1016 0.0923 0.1019 0.1050 0.1024 0.1060 0.95%
-M@10 0.0273 0.0428 0.0443 0.0382 0.0450 0.0461 0.0449 0.0463 0.43%
-M@20 0.0310 0.0471 0.0487 0.0422 0.0494 0.0508 0.0494 0.0508 -
-M@50 0.0347 0.0513 0.0529 0.0464 0.0536 0.0552 0.0536 0.0552 -
 4.1. Experimental Setup
@@
 The performance of our proposed model is evaluated through experiments conducted on three publicly available datasets, which have previously been utilized as evaluation benchmarks in several classical sequential recommendation models.
-  • MovieLens-1M (harper2015movielens): A dataset containing approximately 1 million user ratings for movies, collected from the MovieLens platform.
-  • Amazon-Beauty (mcauley2015image): The user review dataset collected in the beauty category on the Amazon platform was compiled up to the year 2014.
-  • Kuai Rand (kuairand): Acquired from the recommendation logs of the application Kuaishou, the dataset includes millions of interactions involving items that were randomly displayed.
+  • MovieLens-1M (MovieLes2016): A dataset containing approximately 1 million user ratings for movies, collected from the MovieLens platform.
+  • Amazon-Beauty (Amazon2014): The user review dataset collected in the beauty category on the Amazon platform was compiled up to the year 2014.
+  • Kuai Rand (KuaiRand2022): Acquired from the recommendation logs of the application Kuaishou, the dataset includes millions of interactions involving items that were randomly displayed.
 For each user, we sort their interaction records according to the timestamps, thereby generating an interaction sequence for each individual.
-And then retain only those users and items that are associated with a minimum of five interaction records, in accordance with the methodology established in prior research (kang2018self).
+And then retain only those users and items that are associated with a minimum of five interaction records, in accordance with the methodology established in prior research (SASRec2018).
 The statistics of these datasets are shown in Table 2.
-4.1.2. Baseline
-To demonstrate the superiority of our proposed method, we select a set of representative sequential recommendation models to serve as baselines.
-Notably, even though SSD4Rec (ssd4rec) has not been released as open-source, we implemented a variant termed SSD4Rec*, adhering to the traditional sequential recommendation framework to validate the efficacy of our time-aware approach.
-This implementation is strictly based on the SSD architecture and does not incorporate features such as variable-length sequences and bidirectional SSD methods proposed in SSD4Rec.
-The exclusion of these features is intentional, as they are generally regarded as universal enhancement techniques, particularly in the context of handling variable-length sequences.
-  • Caser (tang2018personalized): A classical sequential recommendation model based on CNN.
-  • GRU4Rec (GRU4Rec): A RNN-based method constructed by Gated Recurrent Units (GRU).
-  • SASRec (kang2018self): The first exploration model of applying Transformer in the field of sequential recommendation.
-  • BERT4Rec (sun2019bert4rec): A bidirectional attention sequential recommendation model following the BERT (bert) model paradigm.
-  • TiSASRec (TiSASRec): A time-aware enhanced sequential recommendation model based on SASRec: Considered as paradigm in time-aware sequential recommendation research.
-  • Mamba4Rec (mamba4rec): A pioneering model that explores the application of the Mamba architecture in the domain of sequential recommendation.
-  • SSD4Rec (ssd4rec): The pioneering sequential recommendation model leveraging the SSD architecture, exploits its inherent advantages over the SSM architecture for high-dimensional modeling, and integrates variable-length sequence training techniques to enhance model performance.
-4.1.3. Evaluation Metrics
-We employ well-established recommendation performance metrics including Hit Ratio (HR@K), Normalized Discounted Cumulative Gain (NDCG@K), and Mean Reciprocal Rank (MRR@K) as the criteria for experimental assessment.
-In calculating these metrics, it is necessary to extract the top K scores from the model’s final predicted ranking.
-We choose K values of 10, 20, and 50 to encompass short, medium, and long prediction lengths.
 Table 2. Statistics of the experimented datasets.
@@
 Sparsity 95.15% 99.93% 99.33%
+4.1.2. Baseline
+To demonstrate the superiority of our proposed method, we select a set of representative sequential recommendation models to serve as baselines.
+Notably, even though SSD4Rec (SSD4Rec2024) has not been released as open-source, we implemented a variant termed SSD4Rec*, adhering to the traditional sequential recommendation framework to validate the efficacy of our time-aware approach.
+This implementation is strictly based on the SSD architecture and does not incorporate features such as variable-length sequences and bidirectional SSD methods proposed in SSD4Rec.
+The exclusion of these features is intentional, as they are generally regarded as universal enhancement techniques, particularly in the context of handling variable-length sequences.
+  • Caser (TangW2018): A classical sequential recommendation model based on CNN.
+  • GRU4Rec (GRU4Rec2016): A RNN-based method constructed by Gated Recurrent Units (GRU).
+  • SASRec (SASRec2018): The first exploration model of applying Transformer in the field of sequential recommendation.
+  • BERT4Rec (BERT4Rec2019): A bidirectional attention sequential recommendation model following the BERT (BERT2019) model paradigm.
+  • TiSASRec (TiSASRec2020): A time-aware enhanced sequential recommendation model based on SASRec: Considered as a paradigm in time-aware sequential recommendation research.
+  • LRURec (LRURec2024): A sequential recommendation model based on Linear Recurrent Units (LRU) (LRU2023) with Parallel Scan (Parallelizing2018) optimization.
+  • Mamba4Rec (Mamba4Rec2024): A pioneering model that explores the application of the Mamba architecture in the domain of sequential recommendation.
+  • SSD4Rec (SSD4Rec2024): The pioneering sequential recommendation model leveraging the SSD architecture, exploits its inherent advantages over the SSM architecture for high-dimensional modeling, and integrates variable-length sequence training techniques to enhance model performance.
+Table 3.
+Sequential recommendation performance on the ML-1M, Beauty, and KuaiRand datasets.
+Here, R@k refers to Recall@K, N@K denotes NDCG@K, and M@K represents MRR@K.
+The best and second-best results are indicated in bold and underlined, respectively.
+”Improv.
+SSD” signifies the performance improvement of TiM4Rec compared to the SSD4Rec* results.
+”Improv.
+Sec” signifies the performance improvement of TiM4Rec compared to the second-best results.
+The asterisk in SSD4Rec* denotes that this is our own replicated version.
+The model enclosed in boxes represents the use of time aware enhancement methods.
+Model Type & Model $\rightarrow$ CNN RNN Transformer LRU SSM SSD Improv. SSD Improv. Sec
+Dataset Metric Caser GRU4Rec SASRec BERT4Rec TiSASRec LRURec Mamba4Rec SSD4Rec* TiM4Rec
+MovieLens-1M R@10 0.2156 0.2985 0.3060 0.2800 0.3147 0.3002 0.3253 0.3199 0.3310 3.47% 1.75%
+R@20 0.3083 0.3937 0.4050 0.3853 0.4250 0.4035 0.4354 0.4240 0.4338 2.31% -0.30%
+R@50 0.4515 0.5397 0.5455 0.5377 0.5608 0.5455 0.5707 0.5649 0.5770 2.14% 1.10%
+N@10 0.1157 0.1700 0.1754 0.1584 0.1834 0.1727 0.1891 0.1841 0.1932 4.94% 2.17%
+N@20 0.1391 0.1941 0.2005 0.1852 0.2106 0.1987 0.2169 0.2104 0.2194 4.28% 1.15%
+N@50 0.1675 0.2232 0.2285 0.2154 0.2376 0.2268 0.2436 0.2384 0.2477 3.90% 1.68%
+M@10 0.0853 0.1307 0.1357 0.1214 0.1424 0.1337 0.1474 0.1425 0.1512 6.11% 2.58%
+M@20 0.0917 0.1373 0.1425 0.1288 0.1498 0.1407 0.1551 0.1498 0.1584 5.23% 2.13%
+M@50 0.0963 0.1420 0.1471 0.1337 0.1542 0.1452 0.1593 0.1542 0.1629 5.64% 2.26%
+Amazon Beauty R@10 0.0402 0.0563 0.0851 0.0352 0.0802 0.0840 0.0838 0.0806 0.0854 5.96% 0.35%
+R@20 0.0618 0.0856 0.1194 0.0563 0.1147 0.1191 0.1185 0.1146 0.1204 5.06% 0.84%
+R@50 0.1090 0.1376 0.1759 0.0999 0.1734 0.1799 0.1802 0.1710 0.1800 5.26% -0.11%
+N@10 0.0199 0.0297 0.0425 0.0172 0.0405 0.0427 0.0435 0.0423 0.0446 5.44% 2.53%
+N@20 0.0253 0.0371 0.0511 0.0225 0.0491 0.0522 0.0522 0.0509 0.0533 4.72% 2.11%
+N@50 0.0346 0.0473 0.0623 0.0311 0.0605 0.0640 0.0644 0.0620 0.0651 5.00% 1.09%
+M@10 0.0138 0.0217 0.0294 0.0117 0.0282 0.0300 0.0311 0.0306 0.0321 4.90% 3.22%
+M@20 0.0153 0.0237 0.0317 0.0132 0.0306 0.0326 0.0335 0.0329 0.0345 4.86% 2.99%
+M@50 0.0167 0.0253 0.0335 0.0145 0.0325 0.0345 0.0354 0.0347 0.0363 4.61% 2.54%
+Kuai Rand R@10 0.0801 0.1020 0.1055 0.0938 0.1057 0.1036 0.1094 0.1055 0.1109 5.12% 1.37%
+R@20 0.1344 0.1659 0.1704 0.1537 0.1710 0.1663 0.1768 0.1717 0.1774 3.32% 0.34%
+R@50 0.2561 0.3017 0.3074 0.2873 0.3060 0.3078 0.3154 0.3088 0.3202 3.69% 1.52%
+N@10 0.0395 0.0564 0.0584 0.0510 0.0590 0.0570 0.0608 0.0588 0.0611 3.91% 0.49%
+N@20 0.0531 0.0724 0.0747 0.0660 0.0753 0.0727 0.0777 0.0754 0.0779 3.32% 0.26%
+N@50 0.0770 0.0911 0.1016 0.0923 0.1019 0.1005 0.1050 0.1024 0.1060 3.52% 0.95%
+M@10 0.0273 0.0428 0.0443 0.0382 0.0450 0.0431 0.0461 0.0449 0.0463 3.12% 0.43%
+M@20 0.0310 0.0471 0.0487 0.0422 0.0494 0.0473 0.0508 0.0494 0.0508 2.83% 0.00%
+M@50 0.0347 0.0513 0.0529 0.0464 0.0536 0.0517 0.0552 0.0536 0.0552 2.99% 0.00%
+4.1.3. Evaluation Metrics
+We employ well-established recommendation performance metrics including Hit Ratio (HR@K), Normalized Discounted Cumulative Gain (NDCG@K), and Mean Reciprocal Rank (MRR@K) as the criteria for experimental assessment.
+In calculating these metrics, it is necessary to extract the top K scores from the model’s final predicted ranking.
+We choose K values of 10, 20, and 50 to encompass short, medium, and long prediction lengths.
 4.1.4. Implementation Details
-Our experimental procedures adhere to the paradigm requirements set by the PyTorch (paszke2019pytorch) and RecBole (recbole[1.2.0]).
+Our experimental procedures adhere to the paradigm requirements set by the PyTorch (Pytorch2019) and RecBole (Recbole-1.2.0).
 All models in the experiment are implemented with a model dimension of 64, which is the conventional choice for sequential recommendation models based on pure ID modeling.
-For models based on the Mamba architecture, such as Mamba4Rec, SSD4Rec* and TiM4Rec, the SSM state factor was consistently set to 32, the kernel size for 1D causal convolution is 4, and the block expansion factor for linear projections is 2.
+For models based on the Mamba architecture, such as Mamba4Rec, SSD4Rec*, and TiM4Rec, the SSM state factor was consistently set to 32, the kernel size for 1D causal convolution was 4, and the block expansion factor for linear projections was 2.
 It is notable that the 1D causal convolutional layer within the Mamba architecture constrains the multi-head parameter to multiples of 8.
-However, in the context of sequential recommendation, this parameter typically holds values of 2 or 4.
+However, in the context of sequential recommendation, this parameter typically holds values of 2 or 4 (SASRec2018; BERT4Rec2019).
 Consequently, we modify the 1D causal convolutional layer to accommodate the specific modeling tasks of sequential recommendation, ultimately standardizing it to a value of 4.
@@
 To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M and KuaiRand.
-The Adam optimizer (kingma2014adam) is used uniformly with a learning rate set to 0.01.
+The Adam optimizer (Adam2015) is used uniformly with a learning rate set to 0.01.
 A batch size of 2048 is employed for training, and the validated batch size is 4096.
@@
 However, when compared with models founded on Transformer architectures, it exhibits enhanced performance.
-This observation aligns with our theoretical analysis presented in Section 3.3.
+This observation aligns with our theoretical analysis presented in Section 3.3.1.
   • TiM4Rec exhibits a significant performance enhancement when compared to SSD4Rec*, which is modeled on a purely SSD-based architecture, thereby confirming the efficacy of the Time-aware SSD Block.
 Furthermore, TiM4Rec demonstrates a slight performance advantage over Mamba4Rec.
-This observation underscores the effectiveness of leveraging temporal differences in user interactions to capture shifts in user interests, effectively mitigating the performance decline observed when transitioning from the SSM architecture to the SSD architecture.
+This observation underscores the effectiveness of leveraging time-difference in user interactions to capture shifts in user interests, effectively mitigating the performance decline observed when transitioning from the SSM architecture to the SSD architecture.
 4.3. Dimension comparison
-To assess the suitability of modeling pure ID sequence recommendation tasks in lower dimensions, we compare the experimental results presented in the SSD4Rec paper at 256 dimensions.
+To assess the suitability of modeling pure ID sequential recommendation tasks in lower dimensions, we compare the experimental results presented in the SSD4Rec paper at 256 dimensions.
 These results are shown in Table 4.
@@
 Notably, the performance of SSD4Rec at 256 dimensions is even inferior to that Mamba4Rec at 64 dimensions.
-These findings provide robust evidence that pure ID sequence recommendation tasks are more effectively modeled in lower-dimensional spaces.
+These findings provide robust evidence that pure ID sequential recommendation tasks are more effectively modeled in lower-dimensional spaces.
 Moreover, the computational overhead associated with 256 dimensions is substantially higher, underscoring the advantages of utilizing our model.
@@
 TiSASRec Attention 1149.57s 1.89s
+LRURec LRU 260.92s 0.77s
 Mamba4Rec SSM 111.76s 0.26s
@@
 It is noteworthy that although we only compare the computational efficiency of our own replicated version of SSD4Rec*, the actual computational efficiency of SSD4Rec* surpasses that of SSD4Rec because SSD4Rec* eliminates the bidirectional SSD processing and the variable length sequence methods originally proposed in SSD4Rec, which come with additional overhead.
+Additionally, we compared the training and inference times of various models across different sequence lengths, as shown in Figure 6.
+The results indicate that TiM4Rec consistently maintains a significant speed advantage across all sequence lengths.
+Compared to the Transformer-based SASRec and the SSM-based Mamba4Rec, TiM4Rec effectively inherits the efficiency advantages of the SSD architecture while achieving superior performance.
+This result also demonstrates that our proposed time-aware enhancement method has linear computational complexity, ensuring that TiM4Rec does not experience efficiency degradation as sequence length increases.
+(a) Training time comparison
+(b) Inference time comparison
+Figure 6.
+Comparison of model training and inference times across different sequence lengths on ML-1M.
+The results for SASRec and LRURec at a sequence length of 400 were obtained by reducing the batch size due to Out-Of-Memory.
 4.5. Ablation Study
-To validate the efficacy of our proposed time-aware algorithm along with other critical design elements, we perform a series of ablation studies using the ML-1M dataset.
+To validate the efficacy of our proposed time-aware algorithm along with other critical design elements, we perform a series of ablation studies using the ML-1M and Amazon-Beauty dataset.
 The procedure for the systematic elimination of each component is detailed as follows:
@@
   • N layer : Performance of the Model with N Stacked Layers of Time-aware SSD Blocks.
-Table 6. Results of ablation studies on the ML-1M dataset
-Component R@10 R@20 N@10 N@20 M@10 M@20
-Full 0.3310 0.4338 0.1932 0.2194 0.1512 0.1584
-w/o Time 0.3199 0.4240 0.1841 0.2104 0.1425 0.1498
-w/o FFN 0.3247 0.4278 0.1842 0.2103 0.1415 0.1487
-1 layer 0.3088 0.4146 0.1773 0.2040 0.1373 0.1446
-3 layer 0.3310 0.4323 0.1916 0.2172 0.1490 0.1560
 Based on the results presented in Table 6, it is evident that the time-aware enhancement method our propose significantly improves performance.
 The ablation experiment findings further underscore the critical role of the FFN layer’s feature mapping ability in the extraction of sequence features.
+Table 6. Results of ablation studies
+Dataset $\rightarrow$ ML-1M Amazon-Beauty
+Component R@10 R@20 N@10 N@20 M@10 M@20 R@10 R@20 N@10 N@20 M@10 M@20
+Full 0.3310 0.4338 0.1932 0.2194 0.1512 0.1584 0.0854 0.1204 0.0446 0.0533 0.0321 0.0345
+w/o Time 0.3199 0.4240 0.1841 0.2104 0.1425 0.1498 0.0806 0.1146 0.0423 0.0509 0.0306 0.0329
+w/o FFN 0.3247 0.4278 0.1842 0.2103 0.1415 0.1487 0.0827 0.1183 0.0422 0.0512 0.0298 0.0322
+1 layer 0.3088 0.4146 0.1773 0.2040 0.1373 0.1446 0.0739 0.1042 0.0437 0.0513 0.0344 0.0365
+3 layer 0.3310 0.4323 0.1916 0.2172 0.1490 0.1560 0.0848 0.1211 0.0431 0.0522 0.0303 0.0328
 5. Conclusion and Future Work